### v1

# EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery

Abstract
Recent advances in the visual-language area have developed natural multi-modal large language models (MLLMs) for spatial reasoning through visual prompting. However, due to remote sensing (RS) imagery containing abundant geospatial information that differs from natural images, it is challenging to effectively adapt natural spatial models to the RS domain. Moreover, current RS MLLMs are limited in overly narrow interpretation levels and interaction manner, hindering their applicability in real-world scenarios. To address those challenges, a spatial MLLM named EarthGPT-X is proposed, enabling a comprehensive understanding of multi-source RS imagery, such as optical, synthetic aperture radar (SAR), and infrared. EarthGPT-X offers zoom-in and zoom-out insight, and possesses flexible multi-grained interactive abilities. Moreover, EarthGPT-X unifies two types of critical spatial tasks (i.e., referring and grounding) into a visual prompting framework. To achieve these versatile capabilities, several key strategies are developed. The first is the multi-modal content integration method, which enhances the interplay between images, visual prompts, and text instructions. Subsequently, a cross-domain one-stage fusion training strategy is proposed, utilizing the large language model (LLM) as a unified interface for multi-source multi-task learning. Furthermore, by incorporating a pixel perception module, the referring and grounding tasks are seamlessly unified within a single framework. In addition, the experiments conducted demonstrate the superiority of the proposed EarthGPT-X in multi-grained tasks and its impressive flexibility in multi-modal interaction, revealing significant advancements of MLLM in the RS field.
Index Terms:
Multi-source, Spatial reasoning, Remote sensing (RS), Multi-modal large language models (MLLMs).
I Introduction
Typical spatial understanding tasks in multi-modal learning include referring and grounding, which essentially require aligning regions in the image with text semantics[1]. Referring aims to accurately generate the interpretation sequence of the specific areas. Grounding requires localizing the objects or rendering object segmentation masks based on the user’s language expression[2, 3]. Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focal of research in the natural scenarios[4, 5, 6]. However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[7]. Although current RS MLLMs can handle a wide range of RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception. They solely rely on text instructions which can lead to ambiguity and limited context. It is indispensable for establishing a comprehensive and spatial-aware MLLM to interpret RS images under flexible interactions.
Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[10]. Inspired by this, universal MLLM and multi-grained prompting learning have been extensively studied in the visual-language learning community[11]. Remarkable advances of MLLMs include close-source Open-AI’s GPT series[12, 13, 14], open-source Meta’s LLaMA series[15, 16], and parameter-efficient LLaMA-Adapter series[17, 18], which demonstrate versatile capabilities of handling multi-modal various visual tasks. On top of that, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[19]. Specifically, representative works comprise promptable segmentation model segment anything (SAM)[20], region-of-interest perception model GPT4RoI [4], refer-and-ground MLLM Ferret[1], etc. However, the aforementioned models are tailored for the natural imagery, it is challenging to directly adapt them to the diverse RS data.
Figure 1: An overview of the proposed flexible and comprehensive spatial understanding framework for RS imagery: (a) Task: various multi-grained referring and grounding tasks; (b) Model: the proposed model consists of an image encoder, a visual prompts encoder, a text tokenizer, a LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is transformed based on existing multi-source RS data including optical, SAR, and infrared.
Limited efforts in the RS domain have been made to enhance spatial understanding. A key example is RSPrompter [21], which is based on SAM and develops an automated prompts generation approach to facilitate interactive segmentation. However, RSPrompter is only designed for instance segmentation. RSVG [2] defines the visual grounding tasks in the RS domain and can locate the referred objects using boxes under natural language expression. Recent notable works, such as EarthGPT[8] and Geochat[9], have realized RS holistic and regional imagery comprehension under the unified visual-language learning framework, facilitating the development of MLLMs in the RS domain. Nevertheless, RSVG, EarthGPT, and Geochat solely support language interaction without visual prompts to realize referring expression comprehension. Notably, EarthMarker[22] is constructed as the first visual prompting learning framework in the RS domain and is capable of region-level and point-level RS imagery interpretation. Regretfully, EarthMarker is mainly oriented towards optical images and lacks support for multi-source imagery and free-form visual prompts, hampering flexible human-AI interaction and comprehensive understanding of RS data. Thus, the study on enabling MLLMs to flexibly interpret multi-source RS images at any granularity remains in its infancy stages.
To resolve the limitations of existing models and draw inspiration from how humans understand multi-modal, multi-grained information through flexible interactions, a spatial MLLM named EarthGPT-X is proposed based on visual prompt learning. EarthGPT-X is designed for a comprehensive understanding of multi-source RS imagery under flexible visual prompts (e.g., points, boxes, circles, and scribbles), and unifying referring and grounding tasks within a free-form region-text alignment architecture. As depicted in Fig. 1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, a construction pipeline of training datasets. Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification). Secondly, the key of the model is to realize multi-modal content integration including images, visual prompts, and text instructions. The multi-modal signals are then fed into the large language models (LLM) for subsequent mixed tuning. Additionally, to unify referring and grounding in a single framework, the pixel perception module is designed to generate binary segmentation masks related to referring objects. Thirdly, the training dataset named M-RSVP is constructed, transforming from existing multi-source datasets, which is uniformly reorganized and relabeled as image-visual prompts-text triples, with a portion re-annotated using GPT-4V[23] to support the generation of longer and more detailed descriptions[22].
In addition, observed that existing synthetic aperture radar (SAR) and infrared datasets are limited to object detection tasks, which are insufficient for developing complex visual reasoning capabilities of MLLMs. To address this shortcoming, the multi-domain data are integrated for generalizing general-domain knowledge into SAR and infrared scenarios. Furthermore, to allow for free-form visual prompts data, a noise augmentation method is proposed to add noise that either partially covers the ground truth or extends beyond it to simulate free-form shapes. Based on the newly constructed training datasets, a one-stage all-in-one training strategy is proposed for mixed tuning, facilitating cross-domain data mutual learning and alleviating the need for manual effort for manually setting during the training.
To further investigate EarthGPT-X’s superior comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs. Impressively, EarthGPT-X can realize multi-source data reasoning accurately and provide a more comprehensive and detailed understanding of RS images. Furthermore, to evaluate the effectiveness of the EarthGPT-X’s structure, we conduct ablation studies to assess the designed components’ contributions to the overall performance. Experimental results demonstrate that the proposed EarthGPT-X not only achieves competitive performances on vision-language tasks, but also creates impressive results on referring and grounding tasks.
Our contributions can be summarized as follows.
  • A RS Multi-source Spatial MLLM, EarthGPT-X. A pioneering MLLM named EarthGPT-X is proposed, specifically designed for a comprehensive spatial understanding of multi-source RS imagery. EarthGPT-X advances existing RS MLLMs by enabling flexible multi-grained imagery interpretation, and offers zoom-in and zoom-out insight, powered by designed free-form visual prompts. Furthermore, EarthGPT-X unifies referring and grounding tasks into a visual prompting learning framework, facilitates an adaptable understanding of RS data, and represents a significant milestone in developing versatile MLLMs on RS scenarios.
  • First Multi-source Visual Instruction Dataset, M-RSVP. A multi-source multi-modal visual-prompt-based instruction dataset named M-RSVP, comprising over 200 k image-visual prompts-text triples, is constructed. M-RSVP fills the gap of multi-source visual prompting data in the RS domain, facilitating a multi-grained understanding of RS imagery across various scenes, diverse sources, and varying imaging conditions.
  • Flexible Multi-modal Content Interaction. A multi-modal content integration mechanism is developed, enhancing the mutual understanding between dense image representations, sparse visual prompt features, and text instruction embeddings. This mechanism leads to a deeper comprehension of the referring regions in holistic images. Moreover, by combining with a noise augmentation method for simulating free-form shapes, this mechanism endows the proposed EarthGPT-X with comprehensive RS imagery understanding under flexible human-AI interaction.
  • Cross-domain Knowledge Integration. A cross-domain one-stage fusion training strategy is proposed, transferring and generalizing universal domain-level knowledge into SAR and infrared scenarios. This approach effectively mitigates the limitations of existing SAR and infrared data interpretation, which are typically restricted to object detection tasks and hinder the potential to perform complex multi-task visual reasoning. Furthermore, the one-stage all-in-one training facilitates cross-domain data mutual perception and reduces the reliance on extensive manual effort.
II Related Work
II-A LLMs and MLLMs in Natural Scenarios
The rapid development of LLMs has significantly advanced natural language processing (NLP), demonstrating remarkable capabilities in diverse contexts. Specifically, GPT-3 [12] has demonstrated the power of large-scale models, containing billions to trillions of parameters. InstructGPT [24] and ChatGPT [13] have further enhanced this by showcasing remarkable fluency and adaptability in various conversational tasks. Impressively, the open-source community has significantly contributed to advancing LLMs, with resources like LLaMA series [25] enhancing the instruction-following capabilities of these models. Recent efforts, including Alpaca [26], Vicuna [27], and GPT-4-LLM [28], have introduced full fine-tuning techniques to effectively equip LLMs with robust instruction-following abilities. Additionally, LLaMA-Adapter [18] has demonstrated that parameter-efficient fine-tuning can achieve comparable results with full parameter fine-tuning, offering a more efficient approach to model adaptation.
Notably, an essential step toward achieving general intelligence is integrating multi-modal perception into LLMs, allowing them to handle multi-modal data beyond text. Several works [29, 30, 31, 32] focus on incorporating multi-modal inputs to support task-specific objectives across various modalities. Models like VisualGPT [29], BLIP [30], Flamingo [32], and Kosmos-1 [33] have shown strong multi-modal reasoning abilities by aligning LLMs with visual data. This alignment is achieved using techniques like projection layers, zero-shot attention mechanisms, and intermediate networks, as demonstrated by MiniGPT-4 [34], LLAMA-Adapter V2 [17], and mPLUG-Owl [35]. In addition, with multi-modal LLMs continuing to evolve, other modalities include audio, video [36, 37], and 3D point clouds [38, 39, 40] are being integrated into MLLMs.
II-B MLLMs in RS Scenarios
Inspired by MLLMs in natural domains, MLLMs in the RS field have also made significant strides. Several pioneering RS MLLMs have emerged recently. For example, RSGPT [41], which fine-tunes Instruct-BLIP [42] on a high-quality image-text pairs dataset, demonstrates strong capabilities in image captioning and visual question answer (VQA). However, RSGPT struggles with other visual tasks including object classification and detection, lacking generalization abilities. A key advance is Geochat [9], which is the first versatile MLLM designed to solve multiple tasks for optical RS images. EarthGPT [8] introduces a universal MLLM capable of handling multi-source RS imagery and a broad range of RS visual tasks, and supports regional visual reasoning. In addition, another RS MLLM called LHRS-Bot[43] has established the LHRS-Bench to assist the RS community in evaluating RS-domain MLLMs across diverse evaluation metrics. These models have undoubtedly advanced the development of MLLMs within the RS-specific domain. Nevertheless, EarthGPT, Geochat, and LHRS-Bot solely allow language interaction in multi-turn conversations to handle visual tasks, and fail to support referring comprehension under visual prompt instruction, which restricts flexible interaction in real-world applications.
II-C Prompting Learning
Prompt learning has emerged as a key area of research in NLP [12]. Notable works such as AutoPrompt [44] and CoOp [45] focus on automating prompt template generation for language and vision-language models, reducing the need for manual crafting. In parallel, language prompting has been applied to open-vocabulary detection models, including DetPro [46] and PromptDet [47]. While language prompting is well-developed, visual prompting remains underexplored. A significant advancement in this area is the SAM [20] model, which supports various segmentation prompts to improve zero-shot performance. However, SAM lacks object semantic labels, which is addressed by Semantic-SAM [48]. Moreover, models like GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition. Colorful Prompting Tuning (CPT) [49] uses color-based prompts to enhance MLLMs performance. Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding. Ferret [1] is capable of understanding spatial referring of any shape or granularity within an image and accurately grounding open-vocabulary descriptions. Additionally, SPHINX-V [50] supports free-form visual prompts to achieve pixel-level image comprehension. Mentioned above models are designed for the natural domain.
Note that EarthMarker [22] stands as the pioneering visual prompting framework tailored for the RS domain, specifically designed to interpret RS imagery at both region and point levels. EarthMarker constructs a large-scale RS visual prompting dataset and leverages visual prompts learning to endow MLLM with region-level and point-level RS imagery interpretation capabilities. However, EarthMarker is primarily focused on optical images and lacks support for multi-source imagery and free-form visual prompts. This limitation significantly hinders its potential for flexible interactions and restricts its ability to provide a comprehensive understanding of the diverse and complex characteristics of RS data and broader applicability in real-world RS scenarios.
In summary, existing RS MLLMs have constructed the visual-language learning diagram to comprehend RS imagery and support language interaction in multi-turn dialogues. More advanced than existing MLLMs and EarthMarker, the proposed EarthGPT-X is designed for multi-source imagery comprehension using a dual-prompt system of text instructions and visual prompts. With free-form visual prompting, EarthGPT-X offers comprehensive and flexible visual perception. A straightforward side-by-side comparison of EarthGPT-X versus the recent RS MLLMs is shown in Tab. I.
TABLE I: Comparisons of EarthGPT-X v.s. Recent MLLMs in RS domain.
Model Text Image Visual Prompts Pixel-level
Instruction Optical Multi-source Point Box Free-form Grounding
RSGPT[41]
RS-LLaVA[51]
GeoChat[9]
EarthGPT[8]
Popeye[52]
SkyEyeGPT[53]
EarthMarker[22]
EarthGPT-X
III Methodology
The goal is to unify the referring comprehension at any granularity and pixel grounding into a visual prompt framework in the multi-source RS scenarios. To this end, EarthGPT-X is developed to explore a more general and comprehensive visual prompting learning paradigm. The entire EarthGPT-X framework is detailed in this section. Specifically, we first elaborate on the multi-modal content integration mechanism (Section III-A). Subsequently, the cross-domain one-stage fusion training strategy is presented (Section III-B). In addition, the pixel perception module is introduced in Section III-C.
III-A Multi-modal Content Integration
Figure 2: Overview of the proposed EarthGPT-X model architecture.
The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module. Those multi-modal content encoding and integration are detailed as follows.
III-A1 Visual Global and Local Perception
As shown in Fig. 2 (a), to enhance visual comprehension, the high-resolution inputted images are divided into sub-images for local visual perception. Concurrently, the images are downsampled into lower resolutions and combined with the sub-images for subsequent mixed encoding. The Mixed Visual Encoders (MVE) are designed to encode inputted images and their variants $I^{i}$ into visual representations $V_{\mathrm{img}}$. To leverage the advantages of different network architectures (CNN and Vision Transformer (ViT) [54]) and integrate the local information and contextual dependencies, the semantically complementary backbones CLIP-ConvNeXt [55] and ViT self-supervised by DINOv2[56] are employed as MVE. This visual encoding process can be formulated as
$V_{\mathrm{img}}=\mathrm{Concat}~{}(\mathrm{MVE}~{}(I^{i})),~{}i=1,2,...,N.$ (1)
III-A2 Free-form Spatial-aware
Free-form visual prompts are expected to be flexibly used to guide the model in generating appropriate responses. The shapes of visual prompts include points, boxes, circles, scribbles, etc. To meet the need for flexible interaction at any granularity, a noise augmentation method is developed. Specifically, Gaussian distribution based on the dimensions of the box prompts is sampled to simulate the free-formed shapes. The introduced noise encompasses only a portion of the ground truth or expands it. Subsequently, considering that the visual prompts exhibit varying degrees of sparsity across different shapes, as illustrated in Fig. 2 (b), a randomly initialized ViT is employed as a spatial-aware encoder (SAE) specifically for feature extraction, without any pre-trained weights suitable for image features. The various types of visual prompts $V_{\mathrm{p}}$ are encoded using the designed SAE. The implementation equation can be expressed as follows
$E_{\mathrm{prompt}}=\mathrm{SAE}~{}(V_{\mathrm{p}}),$ (2)
where $E_{\mathrm{prompt}}$ denotes the encoded visual prompts features with free-form shapes.
III-A3 Text Instructions Tokenizer
As displayed in Fig. 2 (c), To facilitate linguistic comprehension, the text instructions $T_{\mathrm{orig}}$ are encoded by the tokenizer[57] module, and projected into language embeddings $L_{\mathrm{instruct}}$, preparing for the subsequent LLM to capture the nuanced interdependence. The detailed process after disassembly is first to use the subword tokenization[57] method to divide text instructions into subwords, enabling understanding of the structure of the linguistic signals. Subsequently, the Word2Vec[58] is adopted to map each subword to a one-dimensional vector embedding. The tokenizer process can be simply described as
$L_{\mathrm{instruct}}=\mathrm{Tokenizer}(~{}T_{\mathrm{orig}}).$ (3)
III-A4 Hybrid Signals Mutual Understanding
To bridge the gap and enhance the interplay between dense visual representations and sparse visual prompts features, a hybrid signal mutual understanding mechanism is proposed. Specifically, as displayed in Fig. 2 (d), the image representations $V_{\mathrm{img}}$ and visual prompts features $E_{\mathrm{prompt}}$ are processed by a self-attention layer and feed-forward layer to obtain $V_{\mathrm{sa}}$ and $E_{\mathrm{ff}}$, respectively. The subsequent cross-attention and second self-attention blocks are used to compute the correlation $H_{\mathrm{ca}}$ between hybrid referring information and multi-scale holistic signals. In essence, this designed mechanism enhances detailed and contextual visual perception, leading to a deeper comprehension of the referring areas in holistic images. The whole procedure can be mathematically represented as
$V_{\mathrm{sa}}$ $=\mathrm{Self\text{-}attention}~{}(V_{\mathrm{img}}),$ (4)
$E_{\mathrm{ff}}$ $=\mathrm{Feed\text{-}forward}~{}(E_{\mathrm{prompt}}),$ (5)
$H_{\mathrm{ca}}$ $=\mathrm{Cross\text{-}attention}~{}(V_{\mathrm{sa}},~{}E_{\mathrm{ff}}),$ (6)
$[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~{}(V_{\mathrm{ca}}+R_{\mathrm{ff}}),$ (7)
where $[V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts. Then, the hybrid visual features are projected to the same dimension as text embeddings through the vision-to-language modality-align projection layer $\mathcal{F}_{\mathrm{vl\text{-}p}}$. This process can be formulated as
$[V\text{-}P]_{\mathrm{proj}}=\mathcal{F}_{\mathrm{vl\text{-}p}}~{}([V\text{-}P]_{\mathrm{hybrid}}).$ (8)
where $[V\text{-}P]_{\mathrm{proj}}$ involve both multi-scale visual signals and free-form spatial-aware information. Notably, the cross-attention layer is updatable during the training.
III-B Cross-domain One-Stage Fusion Training
The previous training pipeline of MLLM routinely uses a step-wise training strategy[8, 22]. However, in the previous pipeline, extensive manual intervention is required. For example, the researcher needs to manually set the different trainable parameters and datasets across various stages. To alleviate the need for manual efforts without sacrificing performance, the cross-domain one-stage fusion training[11] strategy is designed. Most importantly, this strategy further facilitates multi-domain (i.e., natural, optical, SAR, and infrared) data mutual understanding.
In this stage, the projected hybrid tokens are concatenated with the text instruction embeddings $L_{\mathrm{instruct}}$ to develop the multi-modal LLM input $\mathcal{X}$. Subsequently, the integrated multi-modal input is injected into LLM for alignment and deeply interactive understanding (see Fig. 2 (e)). The process can be expressed as
$Y=\mathrm{LLM}~{}(~{}[V\text{-}P]_{\mathrm{proj}},~{}L_{\mathrm{instruct}}).$ (9)
Unlike other models that introduce additional learnable parameters, to simplify the model structure, the Transformer’s attention layers are unfrozen during the training. Specifically, the self-attention head consists of key $K$, query $Q$, and the value $V$, which are implemented by linear layers. The implementations can be expressed as follows
$Q(\mathcal{X})=W_{\mathrm{q}}~{}\mathcal{X}+{b_{\mathrm{q}}},$ (10)
$K(\mathcal{X})=W_{\mathrm{k}}~{}\mathcal{X}+{b_{\mathrm{k}}},$ (11)
$V(\mathcal{X})=W_{\mathrm{v}}~{}\mathcal{X}+{b_{\mathrm{v}}},$ (12)
where $\mathcal{X}$ denotes multi-modal input. The parameters $W_{\mathrm{q}}$, $W_{\mathrm{k}}$, $W_{\mathrm{v}}$, $b_{\mathrm{q}}$, $b_{\mathrm{k}}$, and $b_{\mathrm{v}}$ are updatable during the training.
To summarize, the essence of this process lies in levering multi-domain knowledge to generalize versatile visual reasoning ability into RS multi-source complex scenarios interpretation. Furthermore, through tuning on unified abundant visual prompt-based instruction data, the proposed EarthGPT-X can accomplish referring comprehension of RS imagery with the user’s free-form visual prompts and text instruction-following.
III-C Pixel Perception
To unify referring and grounding in one framework and achieve pixel-level grounding, the pixel encoder-decoder architecture is employed to render the referring object masks, which are shown in Fig. 2 (f). In particular, the special token $pix$ is adopted to represent the pixel grounding masks as embeddings[59]. The embeddings $L_{\mathrm{pix}}$ are generated from the LLM output corresponding to the $pix$ token. Then, the language-to-pixel projection layer $\mathcal{R}_{\mathrm{lp\text{-}p}}$ transforms the embeddings $L_{\mathrm{pix}}$ into the pixel decoder’s feature space. In addition, the grounding encoder is denoted as $\mathcal{G}$, the pixel decoder $\mathcal{P}$. The binary segmentation mask $M$ generated process can be represented as follows
$M=\mathcal{P}\Big{(}\mathcal{R}_{\mathrm{lp}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big{)}.$ (13)
In summary, based on the designed tripartite architecture and all-in-one training strategy, the proposed EarthGPT-X can be expected to effectively and comprehensively realize cross-modal and multi-source knowledge comprehension. Moreover, EarthGPT-X can handle multi-grained RS visual tasks, and flexibly unifies referring and grounding in one framework.
IV Dataset Construction
The goal is to endow the proposed model with flexible and comprehension interpretation abilities on multi-source imagery. However, existing infrared and SAR datasets are limited to object detection tasks, hindering the development of various advanced reasoning capabilities. To address this limitation and unleash the potential of MLLMs to comprehend diverse RS data, a multi-source visual prompt-based instruction dataset named M-RSVP is constructed. Particularly, the natural and optical RS image data, which involves a wide range of visual tasks, are utilized for LLM mixed tuning. In essence, multi-domain data fusion facilitates generalizing the universal domain reasoning capabilities into limited SAR and infrared scenarios. The specific M-RSVP construction pipeline is detailed as follows.
IV-A Raw Data Selection
In order to develop multi-source image understanding abilities of the proposed EarthGPT-X, it is indispensable to utilize data from different scenes and sensors, including outdoor natural images, RS optical images, SAR, and infrared images, covering different visual task types. Thus, the natural scene dataset we selected is OpenImages[60] which meets this characteristic and is ideal for training the model’s visual perception capabilities in general environments. Moreover, we also choose Visual Genome[61] for training, to practice reasoning over multi-region and multi-object interactions. Meanwhile, for the RS optical dataset, we use RSVP[22] which is the first RS visual prompting dataset, allowing the model to develop region/point interpretation skills in diverse geographic and environmental settings. Furthermore, SAR detection datasets includes SSDD[62], AIR-SARShip-2.0 [63], SARDet [64], MSAR [65], and SAR-Aircraft [66]. Infrared object detection contains Sea-shipping, Infrared-security [67], Aerial-mancar [68], Double-light-vehicle [69], and HIT-UAV [70]. Leveraging multi-domain datasets is beneficial for equipping the proposed model to obtain robust generalization capabilities and to handle a wide range of visual tasks. In addition, multi-source data integration ensures comprehensive and flexible image understanding across multiple domains and sensor types.
IV-B Multi-source Data Unification
[…]
In conclusion, all the various datasets from different domains are combined for mixed training to deepen mutual understanding. The multiple free-form visual prompts are allowed to instruct the model to generate suitable answers, ensuring the flexibility of human-AI interaction. Furthermore, multi-domain data fusion is beneficial to realize multi-source multi-task interpretation by leveraging the diverse features and scenarios presented in supplementary image types.
V Experiment
V-A Experimental Setups
V-A1 Implementation Details
The one-stage all-in-one training is conducted in this paper. We use transformer-based decoder-only Llama 2-13B[16] as the basic LLM. The pre-trained DINOv2-ViT L/14 [56] and CLIP-ConvNeXt [55] are adopted as hybrid image encoder MVE. The training image resolution is $448\times 448$. Subsequently, the randomly initialized ViT model is seated as the spatial-aware encoder to refine sparse features. Then, we employ a pre-trained SAM encoder and decoder[20] as the grounding encoder and pixel decoder. The image resolution required to be reshaped for the segmentation model is $1024\times 1024$. Furthermore, we utilize the simple MLPs as the projection layer to realize multi-modal content dimension alignment.
Throughout the training, to keep training efficiency and preserve the generalization ability of backbone models, we only unfreeze the self-attention layers of LLM and the added cross-attention module for tuning. Other modules including visual encoders, projection layers, etc., are kept frozen. The AdamW optimizer[72] with a weight decay of 0.01 and a learning rate of 2e-5 is utilized. The total training is conducted on 8 NVIDIA A100 GPUs with 80 GB memory.
V-A2 Evaluation Metrics
We adopt Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS)[73] as evaluation metrics on referring object classification tasks. Furthermore, we evaluate the performance of the proposed model on the region captioning tasks based on metrics for assessing language generation quality, including BLEU[74], METEOR[75], ROUGE[76], CIDER[77], and SPICE[78]. Those metrics provide a diverse and comprehensive assessment of the quality of identified object information and generated region captions.
TABLE II: Results on referring object classification of EarthGPT-X and other multiple MLLMs on DIOR-RSVG test set.
Methods Publication Year Formats SS S-IoU
Vanilla MLLMs
GeoChat [9] CVPR 2024 Coor 79.59 68.80
Sphinx [79] Arxiv 2023 Coor 93.72 89.37
EarthGPT [8] TGRS 2024 Coor 94.64 90.16
Natural Visual
Prompting MLLMs
Sphinx-V [50] Arxiv 2024 Box 89.07 81.62
Vip-LLava-7b [80] CVPR 2024 Box 72.56 55.94
Vip-LLava-13b [80] CVPR 2024 Box 74.51 60.53
RS Visual
Prompting MLLM
EarthMarker [22] Arxiv 2024 Point 95.96 93.49
EarthMarker [22] Arxiv 2024 Box 98.37 97.24
EarthGPT-X(Ours) Point 96.02 94.12
EarthGPT-X(Ours) Box 98.60 97.83
EarthGPT-X(Ours) Free-form 98.30 96.56
TABLE III: Results on region captioning of EarthGPT-X and other MLLMs on the OPT-RSVG test set.
Method Publication Year Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDER SPICE
Vanilla MLLMs
Qwen-VL-Chat [9] Arxiv 2023 Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
GeoChat [81] CVPR 2024 Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
Sphinx [79] Arxiv 2023 Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
EarthGPT [8] TGRS 2024 Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
Visual Prompting MLLMs
Sphinx-V [50] Arxiv 2024 Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
Vip-LLava-7b [80] CVPR 2024 Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
Vip-LLava-13b [80] CVPR 2024 Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
GLaMM [80] CVPR 2024 Box 12.50 6.06 3.25 1.74 7.98 22.01 23.85 12.40
EarthGPT-X(Ours) Point 52.22 48.13 45.15 42.88 35.01 64.31 477.40 63.14
EarthGPT-X(Ours) Box 64.55 59.76 56.10 53.11 39.22 66.64 487.44 64.20
EarthGPT-X(Ours) Free-form 63.25 58.16 54.89 52.01 38.35 65.01 482.26 63.23
Figure 3: The referring object classification results on multi-source RS images demonstrate the superior region-level RS visual understanding capability of EarthGPT-X, compared to other vanilla and visual prompting MLLMs (the black text indicates consistency with the ground truth, the red text represents incorrect answers).
Figure 4: Example visualization of EarthGPT-X for multi-source (optical, SAR, and infrared) image comprehension and multiple visual tasks reasoning under various visual prompts.
V-B Referring Object Classification
To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla MLLMs and visual prompting MLLMs on DIOR-RSVG[7]. For referring object classification tasks, the text instruction is “Please identify the object category of the marked region in the image”. The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions. The vanilla MLLMs including natural MLLMs Sphinx [79], RS MLLMs GeoChat [81], and EarthGPT [8]. We select the most recent natural visual prompting MLLMs including Sphinx-V [50], Vip-LLava [80], and the RS visual prompting MLLM model EarthMarker[22] as the comparison.
[…]
Figure 5: Pixel grounding examples of EarthGPT-X on multi-source imagery (EarthGPT-X’s interpretation of the object classes is also displayed on the figures).
V-C Region Captioning
For the brief region captioning, the test set of OPT-RSVG [82] is employed for zero-shot evaluation. We elaborately choose both natural and RS-specific MLLMs for performance comparison. Natural scene MLLMs contains Qwen-VL-Chat [9], Sphinx [79], Sphinx-V [50], Vip-LLava [80] and GLaMM [80]. Popular RS-specific MLLMs GeoChat [81] and EarthGPT [8] are selected. The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”. As displayed in Tab. III, the proposed EarthGPT-X significantly outperforms the previous works on all of the evaluation metrics. Notably, using boxes as visual prompts can yield the best performance on region captioning and dialogue ability.
V-D Comprehensive Capabilities Evaluation
Traditional vision models in the RS field primarily focus on a single visual modality and single tasks, exhibiting poor generalization[83, 84, 85]. Thanks to our M-RSVP’s rich multi-sensor knowledge, EarthGPT-X has robust generalization capabilities across optical, SAR, and infrared visual modalities. To verify and assess the comprehensive capabilities of the proposed model, we conduct a thorough qualitative evaluation. Our assessment also covers multi-source images and multiple levels of image interpretation granularity, including image, point, box, and free-form prompted levels. Moreover, the evaluated visual tasks are extensive, such as brief captioning, detailed captioning, ground captioning, region captioning, referring object classification, and inter-relationship analyses. As displayed in Fig. 4, our model demonstrates remarkable visual reasoning and analytical abilities. EarthGPT-X can provide detailed descriptions for multi-source images, as shown in the interpretation results of the first image, which includes overall image depiction, number of targets, target orientation, and scene reasoning. Notably, in the relationship analysis of the second and third images, EarthGPT-X provides detailed and comprehensive responses. Specifically, EarthGPT-X adeptly summarizes the connections between various regions and the overall image relationship, and then analyzes the characteristics and attributes of the targets within each specific area. In summary, EarthGPT-X demonstrates a profound ability to dissect and understand complex visual information in the multi-source RS MLLMs.
V-E Pixel Grounding
To evaluate the pixel grounding capabilities of the proposed model, we specifically chose images that feature complex and different scenes with multiple targets. This experiment aims to test the model’s robustness and precision in identifying and segmenting diverse elements within a single frame. As shown in Fig. 5, the first test image is the port scene, we mark a ship and a tiny vehicle to assess the model’s ability to discern and accurately interpret tiny targets amidst a busy background. Surprisingly, EarthGPT-X not only recognized these smaller objects but also effectively segmented them from the surrounding environment. Another test involved an image from an urban sports center where we marked five distinct targets scattered across different areas, including varied elements such as a swimming pool, tennis court, vehicle, etc. This scenario tests the model’s capacity to handle spatially separated objects within a complex urban setting. EarthGPT-X demonstrates exceptional performances in precise object recognition and pixel-grounding. Moreover, these tests highlight EarthGPT-X’s advanced pixel understanding capabilities and its potential to enhance RS image analyses across different visual modalities.
V-F Ablation Study
In this section, we analyze the impact of key components of EarthGPT-X in detail.
TABLE IV: Ablation results on different visual encoders.
CLIP CLIP DINOv2 SS S-IoU
ConvNeXt-L RN50$\times$16 ViT-L/14
[…]
TABLE V: Ablation results on Image Pre-processing.
Settings Visual Encoder SS S-IoU
Single-scale MVE 94.93 95.05
Multi-scale MVE 96.18 95.29
Multi-scale + Sub-images MVE 98.40 97.13
V-F1 Different Visual Encoders
To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders. Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison. The results are shown in Tab. IV, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieve the most superior performance. The performance of both SS and S-IoU increases accordingly. These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
V-F2 Sub-images Decomposition and Multi-scale
In the visual feature extraction phase of our approach, we enhance the integration of visual features by incorporating multiple scales and high-resolution sub-images, enabling a more profound comprehension of RS images. To evaluate the effectiveness of our modified feature extraction strategy, we conduct ablation studies focusing on the impact of multi-scale features and high-resolution sub-images on the performance of our EarthGPT-X model. As shown in Tab. V, utilizing the multi-scale and sub-images features setting achieves superior performance over single-scale features. These results demonstrate that the detailed analysis of these high-resolution sub-images and more scales allows for capturing finer details that are crucial for accurate interpretation and analysis.
VI Conclusion
In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources. As the first RS spatial model to unify referring and grounding tasks within a visual prompting framework, EarthGPT-X represents a significant advancement in the application of MLLMs to the RS domain. Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery. Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios, such as SAR and infrared. Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-grained visual understanding at the image, region, pixel levels, etc., allowing for precise and intelligent analyses in real-world applications. In the future, we plan to enrich the SAR and infrared datasets to encompass a broader range of tasks, addressing the current limitation of narrow scope. Additionally, we aim to enable EarthGPT-X to perform more complex visual reasoning tasks across multiple visual modalities, further solidifying EarthGPT-X as a versatile and powerful tool for comprehensive RS imagery analyses.

### v2

@@
-EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery
+EarthGPT-X: A Spatial MLLM for Multi-level Multi-Source Remote Sensing Imagery Understanding with Visual Prompting
 Abstract
-Recent advances in the visual-language area have developed natural multi-modal large language models (MLLMs) for spatial reasoning through visual prompting.
-However, due to remote sensing (RS) imagery containing abundant geospatial information that differs from natural images, it is challenging to effectively adapt natural spatial models to the RS domain.
-Moreover, current RS MLLMs are limited in overly narrow interpretation levels and interaction manner, hindering their applicability in real-world scenarios.
-To address those challenges, a spatial MLLM named EarthGPT-X is proposed, enabling a comprehensive understanding of multi-source RS imagery, such as optical, synthetic aperture radar (SAR), and infrared.
-EarthGPT-X offers zoom-in and zoom-out insight, and possesses flexible multi-grained interactive abilities.
-Moreover, EarthGPT-X unifies two types of critical spatial tasks (i.e., referring and grounding) into a visual prompting framework.
-To achieve these versatile capabilities, several key strategies are developed.
-The first is the multi-modal content integration method, which enhances the interplay between images, visual prompts, and text instructions.
-Subsequently, a cross-domain one-stage fusion training strategy is proposed, utilizing the large language model (LLM) as a unified interface for multi-source multi-task learning.
-Furthermore, by incorporating a pixel perception module, the referring and grounding tasks are seamlessly unified within a single framework.
-In addition, the experiments conducted demonstrate the superiority of the proposed EarthGPT-X in multi-grained tasks and its impressive flexibility in multi-modal interaction, revealing significant advancements of MLLM in the RS field.
-Index Terms:
-Multi-source, Spatial reasoning, Remote sensing (RS), Multi-modal large language models (MLLMs).
+Recent advances in natural-domain multi-modal large language models (MLLMs) have demonstrated effective spatial reasoning through visual and textual prompting.
+However, their direct transfer to remote sensing (RS) is hindered by heterogeneous sensing physics, diverse modalities, and unique spatial scales.
+Existing RS MLLMs are mainly limited to optical imagery and plain language interaction, preventing flexible and scalable real-world applications.
+In this article, EarthGPT-X is proposed, the first flexible spatial MLLM that unifies multi-source RS imagery comprehension and accomplishes both coarse-grained and fine-grained visual tasks under diverse visual prompts in a single framework.
+Distinct from prior models, EarthGPT-X introduces: 1) a dual-prompt mechanism combining text instructions with various visual prompts (i.e., point, box, and free-form) to mimic the versatility of referring in human life; 2) a comprehensive multi-source multi-level prompting dataset, the model advances beyond holistic image understanding to support hierarchical spatial reasoning, including scene-level understanding and fine-grained object attributes and relational analysis; 3) a cross-domain one-stage fusion training strategy, enabling efficient and consistent alignment across modalities and tasks.
+Extensive experiments demonstrate that EarthGPT-X substantially outperforms prior nature and RS MLLMs, establishing the first framework capable of multi-source, multi-task, and multi-level interpretation using visual prompting in RS scenarios.
+The code and dataset are available at https://github.com/wivizhang/EarthGPT-X.
 I Introduction
-Typical spatial understanding tasks in multi-modal learning include referring and grounding, which essentially require aligning regions in the image with text semantics[1].
-Referring aims to accurately generate the interpretation sequence of the specific areas.
-Grounding requires localizing the objects or rendering object segmentation masks based on the user’s language expression[2, 3].
-Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focal of research in the natural scenarios[4, 5, 6].
+Typical spatial understanding tasks in multi-modal learning require aligning image regions with text semantics[1].
+Beyond holistic description, such tasks involve accurately interpreting specific regions or objects based on user instructions[2, 3].
+Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focus of research in natural scenarios[4, 5, 6].
 However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[7].
-Although current RS MLLMs can handle a wide range of RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception.
-They solely rely on text instructions which can lead to ambiguity and limited context.
-It is indispensable for establishing a comprehensive and spatial-aware MLLM to interpret RS images under flexible interactions.
+Although existing RS MLLMs can handle multiple RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception.
+They solely rely on text instructions, which can lead to ambiguity and limited context.
+It is indispensable to develop a spatial-aware MLLM to interpret RS imagery under flexible interactions.
 Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[10].
-Inspired by this, universal MLLM and multi-grained prompting learning have been extensively studied in the visual-language learning community[11].
-Remarkable advances of MLLMs include close-source Open-AI’s GPT series[12, 13, 14], open-source Meta’s LLaMA series[15, 16], and parameter-efficient LLaMA-Adapter series[17, 18], which demonstrate versatile capabilities of handling multi-modal various visual tasks.
-On top of that, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[19].
-Specifically, representative works comprise promptable segmentation model segment anything (SAM)[20], region-of-interest perception model GPT4RoI [4], refer-and-ground MLLM Ferret[1], etc.
-However, the aforementioned models are tailored for the natural imagery, it is challenging to directly adapt them to the diverse RS data.
-Figure 1: An overview of the proposed flexible and comprehensive spatial understanding framework for RS imagery: (a) Task: various multi-grained referring and grounding tasks; (b) Model: the proposed model consists of an image encoder, a visual prompts encoder, a text tokenizer, a LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is transformed based on existing multi-source RS data including optical, SAR, and infrared.
+Inspired by this, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[11].
+Representative works include the promptable segmentation model Segment Anything (SAM) [12], which enables class-agnostic delineation from visual or text prompts and shows strong generalization as a versatile backbone.
+The region-of-interest model GPT4RoI [4] integrates large language models (LLMs) for localized reasoning, while Ferret [1] focuses on fine-grained grounding, which can accept diverse region inputs, such as points and bounding boxes.
+Additionally, Wu et al.
+[13] provide a review of visual prompting approaches, including visual prompting, prompt generation, and prompt learning.
+However, designed for natural images, these models cannot be directly adapted to the heterogeneous RS domain due to fundamental differences in sensing physics, modalities, and scales.
+Figure 1: Overview of the proposed framework for flexible and comprehensive spatial understanding in RS imagery: (a) Tasks: a wide range of multi-granularity visual tasks; (b) Model: EarthGPT-X integrates an image encoder, a visual prompt encoder, a text tokenizer, an LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is constructed from multi-source RS data (optical, SAR, infrared) and further enriched with GPT-4V–based annotations.
 Limited efforts in the RS domain have been made to enhance spatial understanding.
@@
 However, RSPrompter is only designed for instance segmentation.
-RSVG [2] defines the visual grounding tasks in the RS domain and can locate the referred objects using boxes under natural language expression.
+RSVG [2] defines the visual grounding tasks in the RS domain and can locate the referred objects using boxes under natural language expressions.
 Recent notable works, such as EarthGPT[8] and Geochat[9], have realized RS holistic and regional imagery comprehension under the unified visual-language learning framework, facilitating the development of MLLMs in the RS domain.
-Nevertheless, RSVG, EarthGPT, and Geochat solely support language interaction without visual prompts to realize referring expression comprehension.
-Notably, EarthMarker[22] is constructed as the first visual prompting learning framework in the RS domain and is capable of region-level and point-level RS imagery interpretation.
-Regretfully, EarthMarker is mainly oriented towards optical images and lacks support for multi-source imagery and free-form visual prompts, hampering flexible human-AI interaction and comprehensive understanding of RS data.
-Thus, the study on enabling MLLMs to flexibly interpret multi-source RS images at any granularity remains in its infancy stages.
-To resolve the limitations of existing models and draw inspiration from how humans understand multi-modal, multi-grained information through flexible interactions, a spatial MLLM named EarthGPT-X is proposed based on visual prompt learning.
-EarthGPT-X is designed for a comprehensive understanding of multi-source RS imagery under flexible visual prompts (e.g., points, boxes, circles, and scribbles), and unifying referring and grounding tasks within a free-form region-text alignment architecture.
+Nevertheless, these models solely support language interaction without visual prompts to realize referring expression comprehension.
+Notably, EarthMarker[15] is a visual prompting model in the RS domain, however, it lacks support for multi-source imagery and free-form visual prompts, hampering flexible human-AI interaction and comprehensive understanding of RS data.
+Thus, enabling MLLMs to interpret multi-source RS imagery at any granularity remains in its infancy.
+To overcome these limitations, we propose EarthGPT-X, a spatial MLLM built upon visual prompt learning, tailored for multi-source RS imagery, supporting diverse visual prompt types (e.g., points, boxes, circles, and scribbles), and enabling comprehensive interpretation from overall scene understanding to fine-grained regional analysis, thereby offering more flexible and scalable visual-language interaction than prior approaches.
 As depicted in Fig.
-1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, a construction pipeline of training datasets.
+1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, and a construction pipeline of training datasets.
 Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification).
-Secondly, the key of the model is to realize multi-modal content integration including images, visual prompts, and text instructions.
-The multi-modal signals are then fed into the large language models (LLM) for subsequent mixed tuning.
-Additionally, to unify referring and grounding in a single framework, the pixel perception module is designed to generate binary segmentation masks related to referring objects.
-Thirdly, the training dataset named M-RSVP is constructed, transforming from existing multi-source datasets, which is uniformly reorganized and relabeled as image-visual prompts-text triples, with a portion re-annotated using GPT-4V[23] to support the generation of longer and more detailed descriptions[22].
-In addition, observed that existing synthetic aperture radar (SAR) and infrared datasets are limited to object detection tasks, which are insufficient for developing complex visual reasoning capabilities of MLLMs.
-To address this shortcoming, the multi-domain data are integrated for generalizing general-domain knowledge into SAR and infrared scenarios.
-Furthermore, to allow for free-form visual prompts data, a noise augmentation method is proposed to add noise that either partially covers the ground truth or extends beyond it to simulate free-form shapes.
-Based on the newly constructed training datasets, a one-stage all-in-one training strategy is proposed for mixed tuning, facilitating cross-domain data mutual learning and alleviating the need for manual effort for manually setting during the training.
-To further investigate EarthGPT-X’s superior comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs.
-Impressively, EarthGPT-X can realize multi-source data reasoning accurately and provide a more comprehensive and detailed understanding of RS images.
-Furthermore, to evaluate the effectiveness of the EarthGPT-X’s structure, we conduct ablation studies to assess the designed components’ contributions to the overall performance.
-Experimental results demonstrate that the proposed EarthGPT-X not only achieves competitive performances on vision-language tasks, but also creates impressive results on referring and grounding tasks.
+Secondly, the key of the model is to realize multi-modal content integration, including images, visual prompts, and text instructions.
+The multi-modal signals are then fed into the LLMs for a one-stage all-in-one training.
+Additionally, a pixel perception module is integrated to extend the model with object segmentation masks.
+Most importantly, the training dataset named M-RSVP is constructed, transforming from existing multi-sensor datasets, which are uniformly reorganized and relabeled as image-visual prompts-text triples.
+In addition, a portion of the data re-annotated using GPT-4V[16] to generate longer and more detailed descriptions.
+A side-by-side comparison of EarthGPT-X with recent RS MLLMs is presented in Tab.
+I, highlighting its advancement in modality coverage, dataset quality, prompting flexibility, and fine-grained multi-task reasoning.
+To further validate EarthGPT-X’s comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs.
+The results indicate that EarthGPT-X achieves accurate multi-source reasoning and provides more detailed RS image understanding.
+Overall, EarthGPT-X not only attains competitive performance on vision–language benchmarks, but also demonstrates remarkable advantages in complex object-level reasoning tasks.
 Our contributions can be summarized as follows.
-  • A RS Multi-source Spatial MLLM, EarthGPT-X.
-A pioneering MLLM named EarthGPT-X is proposed, specifically designed for a comprehensive spatial understanding of multi-source RS imagery.
-EarthGPT-X advances existing RS MLLMs by enabling flexible multi-grained imagery interpretation, and offers zoom-in and zoom-out insight, powered by designed free-form visual prompts.
-Furthermore, EarthGPT-X unifies referring and grounding tasks into a visual prompting learning framework, facilitates an adaptable understanding of RS data, and represents a significant milestone in developing versatile MLLMs on RS scenarios.
-  • First Multi-source Visual Instruction Dataset, M-RSVP.
-A multi-source multi-modal visual-prompt-based instruction dataset named M-RSVP, comprising over 200 k image-visual prompts-text triples, is constructed.
-M-RSVP fills the gap of multi-source visual prompting data in the RS domain, facilitating a multi-grained understanding of RS imagery across various scenes, diverse sources, and varying imaging conditions.
-  • Flexible Multi-modal Content Interaction.
-A multi-modal content integration mechanism is developed, enhancing the mutual understanding between dense image representations, sparse visual prompt features, and text instruction embeddings.
-This mechanism leads to a deeper comprehension of the referring regions in holistic images.
-Moreover, by combining with a noise augmentation method for simulating free-form shapes, this mechanism endows the proposed EarthGPT-X with comprehensive RS imagery understanding under flexible human-AI interaction.
-  • Cross-domain Knowledge Integration.
-A cross-domain one-stage fusion training strategy is proposed, transferring and generalizing universal domain-level knowledge into SAR and infrared scenarios.
-This approach effectively mitigates the limitations of existing SAR and infrared data interpretation, which are typically restricted to object detection tasks and hinder the potential to perform complex multi-task visual reasoning.
-Furthermore, the one-stage all-in-one training facilitates cross-domain data mutual perception and reduces the reliance on extensive manual effort.
+  • EarthGPT-X: First Multi-source Spatial MLLM.
+We propose EarthGPT-X, a spatial MLLM designed for multi-source (i.e., optical, SAR, and infrared) RS imagery understanding.
+It advances existing RS MLLMs by supporting free-form visual prompts (e.g., points, boxes, scribbles) to enable flexible multi-granularity interaction.
+EarthGPT-X unifies diverse spatial reasoning tasks into a visual prompting framework, providing a generalizable interpretation across heterogeneous RS data.
+  • M-RSVP: First Multi-source Visual Prompting Dataset.
+The first multi-source visual prompting dataset for RS, titled M-RSVP, consisting of over 650k image-visual prompt-text triples, is constructed.
+The constructed M-RSVP bridges the gap of multi-source visual prompting data in the RS domain, enabling multi-granularity interpretation across diverse scenes and sensors.
+  • Multi-domain Multi-modal Multi-level Interaction.
+A hybrid signal mutual understanding mechanism is designed to strengthen the interplay between dense image features, sparse visual prompts, and text instructions, thereby improving the interpretation of referring regions within holistic images.
+Moreover, our cross-domain one-stage fusion training significantly streamlines the training procedure of MLLMs, while ensuring seamless integration of multi-source knowledge.
+  • Superior Performance.
+Extensive experiments benchmark EarthGPT-X against SOTA natural and RS MLLMs across diverse multi-level visual tasks and multi-source imagery.
+Furthermore, EarthGPT-X also excels in complex reasoning tasks, such as object inter-relationship analysis.
+These advances collectively represent a significant milestone toward developing versatile and flexible MLLMs for RS scenarios.
+TABLE I: Comparisons of EarthGPT-X v.s. Recent MLLMs.
+Model Textual Modalities RS Dataset Tasks Visual Prompts Pix.
+Prompt Opt. Multi. Pub. Refi. Regu. Comp. Point Box Free-form Grd.
+Shikra (Arxiv 2023)[6]
+GPT4RoI (ECCV 2024)[17]
+ViP-LLaVA (CVPR 2024)[18]
+RS-LLaVA (RS 2024)[19]
+GeoChat (CVPR 2024)[9]
+EarthGPT (TGRS 2024)[8]
+Popeye (JSTAR 2024)[20]
+LHRS-Bot (ECCV 2024)[21]
+VHM (AAAI 2025)[22]
+EarthDial (CVPR 2025)[23]
+EarthMarker (TGRS 2025)[15]
+EarthGPT-X (Ours)
+Note: Comparisons in terms of modalities (optical or multi-source), dataset construction (converted directly from public data or Refined using powerful MLLMs such as GPT-4V or Gemini), supported tasks (Regular visual tasks or complex reasoning such as inter-relationship analysis), visual prompting capabilities, and extended pixel-level grounding.
 II Related Work
-II-A LLMs and MLLMs in Natural Scenarios
-The rapid development of LLMs has significantly advanced natural language processing (NLP), demonstrating remarkable capabilities in diverse contexts.
-Specifically, GPT-3 [12] has demonstrated the power of large-scale models, containing billions to trillions of parameters.
-InstructGPT [24] and ChatGPT [13] have further enhanced this by showcasing remarkable fluency and adaptability in various conversational tasks.
-Impressively, the open-source community has significantly contributed to advancing LLMs, with resources like LLaMA series [25] enhancing the instruction-following capabilities of these models.
-Recent efforts, including Alpaca [26], Vicuna [27], and GPT-4-LLM [28], have introduced full fine-tuning techniques to effectively equip LLMs with robust instruction-following abilities.
-Additionally, LLaMA-Adapter [18] has demonstrated that parameter-efficient fine-tuning can achieve comparable results with full parameter fine-tuning, offering a more efficient approach to model adaptation.
-Notably, an essential step toward achieving general intelligence is integrating multi-modal perception into LLMs, allowing them to handle multi-modal data beyond text.
+II-A MLLMs in Natural Scenarios
+The rapid development of LLMs has significantly advanced natural language processing (NLP), demonstrating remarkable capabilities in diverse contexts [24, 25, 26].
+Then, an essential step toward achieving general intelligence is integrating multi-modal perception into LLMs, allowing them to handle multi-modal data beyond text.
 Several works [27, 28, 29, 30] focus on incorporating multi-modal inputs to support task-specific objectives across various modalities.
-Models like VisualGPT [29], BLIP [30], Flamingo [32], and Kosmos-1 [33] have shown strong multi-modal reasoning abilities by aligning LLMs with visual data.
-This alignment is achieved using techniques like projection layers, zero-shot attention mechanisms, and intermediate networks, as demonstrated by MiniGPT-4 [34], LLAMA-Adapter V2 [17], and mPLUG-Owl [35].
-In addition, with multi-modal LLMs continuing to evolve, other modalities include audio, video [36, 37], and 3D point clouds [38, 39, 40] are being integrated into MLLMs.
-II-B MLLMs in RS Scenarios
-Inspired by MLLMs in natural domains, MLLMs in the RS field have also made significant strides.
-Several pioneering RS MLLMs have emerged recently.
-For example, RSGPT [41], which fine-tunes Instruct-BLIP [42] on a high-quality image-text pairs dataset, demonstrates strong capabilities in image captioning and visual question answer (VQA).
-However, RSGPT struggles with other visual tasks including object classification and detection, lacking generalization abilities.
-A key advance is Geochat [9], which is the first versatile MLLM designed to solve multiple tasks for optical RS images.
-EarthGPT [8] introduces a universal MLLM capable of handling multi-source RS imagery and a broad range of RS visual tasks, and supports regional visual reasoning.
-In addition, another RS MLLM called LHRS-Bot[43] has established the LHRS-Bench to assist the RS community in evaluating RS-domain MLLMs across diverse evaluation metrics.
-These models have undoubtedly advanced the development of MLLMs within the RS-specific domain.
-Nevertheless, EarthGPT, Geochat, and LHRS-Bot solely allow language interaction in multi-turn conversations to handle visual tasks, and fail to support referring comprehension under visual prompt instruction, which restricts flexible interaction in real-world applications.
+Models like VisualGPT [27], Flamingo [30], and SPHINX-X[31] have shown strong multi-modal reasoning abilities by aligning LLMs with visual data.
+This alignment is achieved through projection layers, zero-shot attention, and intermediate networks, as in MiniGPT-4 [32], LLAMA-Adapter V2 [33], and mPLUG-Owl [34].
+More recent advances include Qwen2.5-VL [35], MiniCPM-V2.6 [36], InternVL2.5 [37], InternLM-XComposer2.5 [38], and DeepSeek-VL2 [39].
+These enhanced versions advance prior frameworks by broadening input modalities, leveraging higher-quality large-scale datasets, and extending the scope of multi-modal reasoning tasks.
+II-B Specialist and MLLMs in RS Scenarios
+RS imagery encompasses diverse modalities such as SAR, optical, and infrared, each capturing complementary physical characteristics of the observed scene.
+Recent advanced SAR detection adopts generative or predictive paradigms, e.g., MaDiNet [40] with Gamma Diffusion–based bounding-box generation and SAR-JEPA [41] with joint-embedding predictive recognition.
+In optical understanding, specialist models such as CSDS [42] are designed for scene classification, while Qonly [43] targets VQA tasks.
+For infrared detection, IRSAM [44] enhances SAM to achieve more discriminative representations.
+Although effective for their respective tasks, these specialist models lack the ability to perform multiple visual tasks in a unified architecture, thus limiting scalability and hindering broader real-world deployment.
+Inspired by the success of MLLMs in natural image domains, RS research has also begun to adopt MLLMs with notable progress.
+RSGPT [45], built on Instruct-BLIP [46], demonstrates strong performance in captioning and VQA but struggles with classification and detection.
+Cross-modal transfer learning in the RS MLLM field has recently gained traction, with representative works exemplifying different stages of progress [47, 48].
+GeoChat [9] represents the first versatile RS MLLM for multiple optical tasks, while SkyEyeGPT [21] further extends broader task coverage, EarthGPT [8] bridging MLLMs with optical, SAR, and infrared domains.
+LHRS-Bot [49] additionally contributes by introducing LHRS-Bench for standardized evaluation.
+These efforts are collectively paving the way for more unified and versatile frameworks.
+However, they remain limited to text-only prompts, lacking flexible interactive comprehension under visual instructions.
 II-C Prompting Learning
-Prompt learning has emerged as a key area of research in NLP [12].
-Notable works such as AutoPrompt [44] and CoOp [45] focus on automating prompt template generation for language and vision-language models, reducing the need for manual crafting.
-In parallel, language prompting has been applied to open-vocabulary detection models, including DetPro [46] and PromptDet [47].
-While language prompting is well-developed, visual prompting remains underexplored.
-A significant advancement in this area is the SAM [20] model, which supports various segmentation prompts to improve zero-shot performance.
-However, SAM lacks object semantic labels, which is addressed by Semantic-SAM [48].
-Moreover, models like GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition.
-Colorful Prompting Tuning (CPT) [49] uses color-based prompts to enhance MLLMs performance.
+Prompt learning has become a key research direction in NLP [24], with works such as AutoPrompt [50] and CoOp [51] automating prompt design for language and vision–language models.
+Subsequently, the idea evolved into visual prompting [13], marking a broader shift from language-based to multi-modal prompts.
+A milestone in this line is the Segment Anything Model (SAM) [12], which enables class-agnostic segmentation via point, box, or text prompts and demonstrates strong generalization across diverse domains.
+GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition.
+Colorful Prompting Tuning (CPT) [52] uses color-based prompts to enhance MLLMs’ performance.
 Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding.
-Ferret [1] is capable of understanding spatial referring of any shape or granularity within an image and accurately grounding open-vocabulary descriptions.
-Additionally, SPHINX-V [50] supports free-form visual prompts to achieve pixel-level image comprehension.
-Mentioned above models are designed for the natural domain.
-Note that EarthMarker [22] stands as the pioneering visual prompting framework tailored for the RS domain, specifically designed to interpret RS imagery at both region and point levels.
-EarthMarker constructs a large-scale RS visual prompting dataset and leverages visual prompts learning to endow MLLM with region-level and point-level RS imagery interpretation capabilities.
-However, EarthMarker is primarily focused on optical images and lacks support for multi-source imagery and free-form visual prompts.
-This limitation significantly hinders its potential for flexible interactions and restricts its ability to provide a comprehensive understanding of the diverse and complex characteristics of RS data and broader applicability in real-world RS scenarios.
-In summary, existing RS MLLMs have constructed the visual-language learning diagram to comprehend RS imagery and support language interaction in multi-turn dialogues.
-More advanced than existing MLLMs and EarthMarker, the proposed EarthGPT-X is designed for multi-source imagery comprehension using a dual-prompt system of text instructions and visual prompts.
-With free-form visual prompting, EarthGPT-X offers comprehensive and flexible visual perception.
-A straightforward side-by-side comparison of EarthGPT-X versus the recent RS MLLMs is shown in Tab.
-I.
-TABLE I: Comparisons of EarthGPT-X v.s. Recent MLLMs in RS domain.
-Model Text Image Visual Prompts Pixel-level
-Instruction Optical Multi-source Point Box Free-form Grounding
-RSGPT[41]
-RS-LLaVA[51]
-GeoChat[9]
-EarthGPT[8]
-Popeye[52]
-SkyEyeGPT[53]
-EarthMarker[22]
-EarthGPT-X
+ViP-LLava[47] supports arbitrary visual cues like scribbles and arrows to interpret images flexibly.
+Wu et al.
+[13] provide a comprehensive survey of visual prompting approaches, covering visual prompting, prompt generation, and prompt learning.
+These models are primarily designed for the natural image domain and fail to generalize to the RS multi-source data.
+Notably, EarthMarker [15] is a visual prompting framework tailored for the RS data, but it remains confined to optical imagery and does not support flexible prompt types.
+These gaps highlight the necessity of a specialized model that can support multi-source imagery and flexible interaction in RS scenarios, motivating the development of the proposed model.
 III Methodology
-The goal is to unify the referring comprehension at any granularity and pixel grounding into a visual prompt framework in the multi-source RS scenarios.
-To this end, EarthGPT-X is developed to explore a more general and comprehensive visual prompting learning paradigm.
-The entire EarthGPT-X framework is detailed in this section.
-Specifically, we first elaborate on the multi-modal content integration mechanism (Section III-A).
-Subsequently, the cross-domain one-stage fusion training strategy is presented (Section III-B).
-In addition, the pixel perception module is introduced in Section III-C.
+The EarthGPT-X framework is detailed in this section.
+We first elaborate on the multi-modal content integration mechanism (Section III-A).
+Then, the cross-domain one-stage fusion training strategy is presented (Section III-B).
+The extended pixel perception module is introduced in Section III-C.
 III-A Multi-modal Content Integration
-Figure 2: Overview of the proposed EarthGPT-X model architecture.
+Figure 2: Overview of the proposed EarthGPT-X architecture.
+It integrates images, visual prompts, and text instructions through a hybrid signal interaction mechanism, and employs a unified one-stage fusion training strategy to enable multi-source knowledge integration and multi-level multi-task reasoning.
 The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module.
-Those multi-modal content encoding and integration are detailed as follows.
+The multi-modal content encoding and integration are detailed as follows.
 III-A1 Visual Global and Local Perception
-As shown in Fig.
-2 (a), to enhance visual comprehension, the high-resolution inputted images are divided into sub-images for local visual perception.
+As shown at the bottom left of Fig.
+2, to enhance visual comprehension, the high-resolution input images are divided into sub-images for local visual perception.
 Concurrently, the images are downsampled into lower resolutions and combined with the sub-images for subsequent mixed encoding.
-The Mixed Visual Encoders (MVE) are designed to encode inputted images and their variants $I^{i}$ into visual representations $V_{\mathrm{img}}$.
+The Mixed Visual Encoders (MVE) are designed to encode input images and their variants $I^{i}$ into visual representations $V_{\mathrm{img}}$.
 To leverage the advantages of different network architectures (CNN and Vision Transformer (ViT) [53]) and integrate the local information and contextual dependencies, the semantically complementary backbones CLIP-ConvNeXt [54] and ViT self-supervised by DINOv2[55] are employed as MVE.
@@
 The introduced noise encompasses only a portion of the ground truth or expands it.
-Subsequently, considering that the visual prompts exhibit varying degrees of sparsity across different shapes, as illustrated in Fig.
-2 (b), a randomly initialized ViT is employed as a spatial-aware encoder (SAE) specifically for feature extraction, without any pre-trained weights suitable for image features.
+Subsequently, considering that the visual prompts exhibit varying degrees of sparsity across different shapes, as illustrated in the bottom right of Fig.
+2, a randomly initialized ViT is employed as a spatial-aware encoder (SAE) specifically for feature extraction, without any pre-trained weights suitable for image features.
 The various types of visual prompts $V_{\mathrm{p}}$ are encoded using the designed SAE.
@@
 III-A3 Text Instructions Tokenizer
-As displayed in Fig.
-2 (c), To facilitate linguistic comprehension, the text instructions $T_{\mathrm{orig}}$ are encoded by the tokenizer[57] module, and projected into language embeddings $L_{\mathrm{instruct}}$, preparing for the subsequent LLM to capture the nuanced interdependence.
+As displayed in left of Fig.
+2, to facilitate linguistic comprehension, the text instructions $T_{\mathrm{orig}}$ are encoded by the tokenizer[56] module, and projected into language embeddings $L_{\mathrm{instruct}}$, preparing for the subsequent LLM to capture the nuanced interdependence.
 The detailed process after disassembly is first to use the subword tokenization[56] method to divide text instructions into subwords, enabling understanding of the structure of the linguistic signals.
@@
 To bridge the gap and enhance the interplay between dense visual representations and sparse visual prompts features, a hybrid signal mutual understanding mechanism is proposed.
-Specifically, as displayed in Fig.
-2 (d), the image representations $V_{\mathrm{img}}$ and visual prompts features $E_{\mathrm{prompt}}$ are processed by a self-attention layer and feed-forward layer to obtain $V_{\mathrm{sa}}$ and $E_{\mathrm{ff}}$, respectively.
+Specifically, as displayed in the middle of Fig.
+2, the image representations $V_{\mathrm{img}}$ and visual prompts features $E_{\mathrm{prompt}}$ are processed by a self-attention layer and feed-forward layer to obtain $V_{\mathrm{sa}}$ and $E_{\mathrm{ff}}$, respectively.
 The subsequent cross-attention and second self-attention blocks are used to compute the correlation $H_{\mathrm{ca}}$ between hybrid referring information and multi-scale holistic signals.
@@
 $[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~(V_{\mathrm{ca}}+R_{\mathrm{ff}}),$ (7)
-where $[V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts.
+where $V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts.
 Then, the hybrid visual features are projected to the same dimension as text embeddings through the vision-to-language modality-align projection layer $\mathcal{F}_{\mathrm{vl\text{-}p}}$.
 This process can be formulated as
-$[V\text{-}P]_{\mathrm{proj}}=\mathcal{F}_{\mathrm{vl\text{-}p}}~{}([V\text{-}P]_{\mathrm{hybrid}}).$ (8)
-where $[V\text{-}P]_{\mathrm{proj}}$ involve both multi-scale visual signals and free-form spatial-aware information. Notably, the cross-attention layer is updatable during the training.
+$V\text{-}P]_{\mathrm{proj}}=\mathcal{F}_{\mathrm{vl\text{-}p}}~([V\text{-}P]_{\mathrm{hybrid}}).$ (8)
+where $V\text{-}P]_{\mathrm{proj}}$ involve both multi-scale visual signals and free-form spatial-aware information. Notably, the cross-attention layer is updatable during training.
 III-B Cross-domain One-Stage Fusion Training
-The previous training pipeline of MLLM routinely uses a step-wise training strategy[8, 22].
-However, in the previous pipeline, extensive manual intervention is required.
-For example, the researcher needs to manually set the different trainable parameters and datasets across various stages.
-To alleviate the need for manual efforts without sacrificing performance, the cross-domain one-stage fusion training[11] strategy is designed.
-Most importantly, this strategy further facilitates multi-domain (i.e., natural, optical, SAR, and infrared) data mutual understanding.
+The conventional training pipeline of MLLMs typically follows a two-stage approach [8, 15], which requires extensive manual intervention.
+In particular, researchers must carefully configure trainable parameters and dataset combinations across different stages, making the process complex and labor-intensive.
+Different from the prior models, we adopt a cross-domain one-stage fusion training strategy [31], which streamlines the process by uniformly treating all collected datasets while preserving strong performance.
+More importantly, this unified paradigm facilitates mutual understanding across heterogeneous domains, including natural, optical, SAR, and infrared imagery.
 In this stage, the projected hybrid tokens are concatenated with the text instruction embeddings $L_{\mathrm{instruct}}$ to develop the multi-modal LLM input $\mathcal{X}$.
 Subsequently, the integrated multi-modal input is injected into LLM for alignment and deeply interactive understanding (See Fig.
-2 (e)).
+2 LLM part).
 The process can be expressed as
@@
 Unlike other models that introduce additional learnable parameters, to simplify the model structure, the Transformer’s attention layers are unfrozen during the training.
-Specifically, the self-attention head consists of key $K$, query $Q$, and the value $V$, which are implemented by linear layers.
+Specifically, the self-attention head consists of key $K$, query $Q$, and value $V$, which are implemented by linear layers.
 The implementations can be expressed as follows
@@
 The parameters $W_{\mathrm{q}}$, $W_{\mathrm{k}}$, $W_{\mathrm{v}}$, $b_{\mathrm{q}}$, $b_{\mathrm{k}}$, and $b_{\mathrm{v}}$ are updatable during the training.
-To summarize, the essence of this process lies in levering multi-domain knowledge to generalize versatile visual reasoning ability into RS multi-source complex scenarios interpretation.
-Furthermore, through tuning on unified abundant visual prompt-based instruction data, the proposed EarthGPT-X can accomplish referring comprehension of RS imagery with the user’s free-form visual prompts and text instruction-following.
-III-C Pixel Perception
-To unify referring and grounding in one framework and achieve pixel-level grounding, the pixel encoder-decoder architecture is employed to render the referring object masks, which are shown in Fig.
-2 (f).
-In particular, the special token $pix$ is adopted to represent the pixel grounding masks as embeddings[59].
-The embeddings $L_{\mathrm{pix}}$ are generated from the LLM output corresponding to the $pix$ token.
+To summarize, EarthGPT-X leverages multi-domain knowledge to generalize versatile visual reasoning into multi-source RS scenarios.
+By tuning on abundant unified dual-prompting data, it achieves flexible referring comprehension with free-form visual prompts and text instructions.
+III-C Integrated with Pixel Perception Module
+To further enhance spatial perception, we integrate a pixel perception module into EarthGPT-X, enabling pixel-level grounding under flexible visual prompts.
+In particular, the special token pixpix is adopted to represent the pixel grounding masks as embeddings[58].
+The embeddings $L_{\mathrm{pix}}$ are generated from the LLM output corresponding to the pixpix token.
 Then, the language-to-pixel projection layer $\mathcal{R}_{\mathrm{lp\text{-}p}}$ transforms the embeddings $L_{\mathrm{pix}}$ into the pixel decoder’s feature space.
@@
 $M=\mathcal{P}\Big(\mathcal{R}_{\mathrm{lp}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big).$ (13)
-In summary, based on the designed tripartite architecture and all-in-one training strategy, the proposed EarthGPT-X can be expected to effectively and comprehensively realize cross-modal and multi-source knowledge comprehension.
-Moreover, EarthGPT-X can handle multi-grained RS visual tasks, and flexibly unifies referring and grounding in one framework.
+In summary, EarthGPT-X, built on a tripartite architecture and an all-in-one training strategy, enables effective cross-modal and multi-source knowledge comprehension.
+It further supports multi-grained RS visual tasks with flexible integration of referring and grounding in a unified framework.
+TABLE II: Statistics of Training Data.
+Tasks Raw Data #Samples Domain
+Image Captioning RSICD [59] 24.3k Optical
+UCM-Captions [60] 10k
+Scene Classification RSSCN7 [61] 0.56k
+EuroSAT [62] 5.4k
+ROC (point) SIOR [63] 200k
+Region Captioning DIOR-RSVG [2] 31.5k
+ROC (box) AIR-SARShip [64] 1.43k SAR
+SSDD [65] 1.86k
+SARDet [66] 106k
+SAR Aircraft [67] 1.9k
+SAR-Ship [68] 33.7k
+MSAR [69] 32.5k
+ROC (box) Sea-shipping[70] 8.4k Infrared
+Infrared-security [71] 9.6k
+Aerial-mancar [72] 31.6k
+Double-light-vehicle [73] 5.1k
+Captioning,
+Relationship Analysis
+(GPT-assisted)
+AIR-SARShip [64] 0.5k SAR
+SARDet [66] 106k
+SSDD [65] 1.0k
+SAR Aircraft [67] 1.9k
+Infrared-security [71] 9.6k Infrared
+Aerial-mancar [72] 31.6k
+Notes: ‘ROC’ denotes Referring Object Classification. The ‘box’ and ‘point’ refer to box-level and point-level visual prompts, respectively.
+Figure 3: (a) Box prompts: An example of converting SAR object detection data into multi-modal conversation training data.
+The training data consists of three parts: the image, the visual prompt, and the text (question and answer).
+The image is taken directly from the object detection dataset, while the visual prompts are from the ground-truth bounding boxes.
+The text prompts can be selected from fixed prompt templates (*note: enhanced by GPT-4V).
+(b) Point prompts: each image in the semantic segmentation dataset is divided into 32 × 32 patches, randomly sampled within each patch as the point prompts, with the category retrieved from the corresponding segmentation map.
+(c) Free-form prompts: Simulated by applying random noise-based augmentations to regions of interest.
+These free-form cues emulate flexible user interactions, enabling the model to generalize beyond structured visual prompts.
+Figure 4: Illustration of the standardized prompt templates designed for GPT-4V annotation generation.
+The (a) part shows the public templates for generic data construction, while the (b) part demonstrates the customized task templates, $\mathrm{Role}>$ and $\mathrm{Format}>$, tailored to multi-source RS imagery.
+These templates embed object categories explicitly, ensuring the correctness and consistency of generated annotations.
 IV Dataset Construction
-The goal is to endow the proposed model with flexible and comprehension interpretation abilities on multi-source imagery.
-However, existing infrared and SAR datasets are limited to object detection tasks, hindering the development of various advanced reasoning capabilities.
-To address this limitation and unleash the potential of MLLMs to comprehend diverse RS data, a multi-source visual prompt-based instruction dataset named M-RSVP is constructed.
-Particularly, the natural and optical RS image data, which involves a wide range of visual tasks, are utilized for LLM mixed tuning.
-In essence, multi-domain data fusion facilitates generalizing the universal domain reasoning capabilities into limited SAR and infrared scenarios.
-The specific M-RSVP construction pipeline is detailed as follows.
-IV-A Raw Data Selection
-In order to develop multi-source image understanding abilities of the proposed EarthGPT-X, it is indispensable to utilize data from different scenes and sensors, including outdoor natural images, RS optical images, SAR, and infrared images, covering different visual task types.
-Thus, the natural scene dataset we selected is OpenImages[60] which meets this characteristic and is ideal for training the model’s visual perception capabilities in general environments.
-Moreover, we also choose Visual Genome[61] for training, to practice reasoning over multi-region and multi-object interactions.
-Meanwhile, for the RS optical dataset, we use RSVP[22] which is the first RS visual prompting dataset, allowing the model to develop region/point interpretation skills in diverse geographic and environmental settings.
-Furthermore, SAR detection datasets includes SSDD[62], AIR-SARShip-2.0 [63], SARDet [64], MSAR [65], and SAR-Aircraft [66].
-Infrared object detection contains Sea-shipping, Infrared-security [67], Aerial-mancar [68], Double-light-vehicle [69], and HIT-UAV [70].
-Leveraging multi-domain datasets is beneficial for equipping the proposed model to obtain robust generalization capabilities and to handle a wide range of visual tasks.
-In addition, multi-source data integration ensures comprehensive and flexible image understanding across multiple domains and sensor types.
-IV-B Multi-source Data Unification
-[…]
-In conclusion, all the various datasets from different domains are combined for mixed training to deepen mutual understanding.
-The multiple free-form visual prompts are allowed to instruct the model to generate suitable answers, ensuring the flexibility of human-AI interaction.
-Furthermore, multi-domain data fusion is beneficial to realize multi-source multi-task interpretation by leveraging the diverse features and scenarios presented in supplementary image types.
+To support comprehensive multi-source RS understanding, we construct a large-scale dataset called M-RSVP, which comprises over 650k samples spanning various tasks, modalities, and annotation formats.
+The detailed introduction is as follows.
+IV-A Multi-Source Multi-Domain Data Construction
+For developing multi-source and versatile image understanding abilities of EarthGPT-X, it is indispensable to utilize data from different scenes and sensors, including outdoor natural images, RS optical images, SAR, and infrared images, covering different visual task types.
+Thus, as shown in Tab.
+II, we construct a multi-domain collection of raw data covering diverse vision-language tasks.
+For the RS optical modality, we adopt RSVP [15], the first visual prompting dataset in the RS domain, which enables the model to learn region- and point-level understanding in complex geographic environments.
+Meanwhile, we incorporate more optical datasets [63, 2] to enrich the diversity of tasks and enhance the model’s generalization in RS scenarios.
+For SAR and infrared modalities, visual prompting data are mainly derived from publicly available object detection datasets, which are reorganized and annotated in a unified instruction format.
+Additionally, the natural scene dataset we selected is OpenImages [74] and Visual Genome [75], which meets this characteristic and is ideal for training the model’s visual perception capabilities in general environments.
+The multi-domain data integration ensures comprehensive and flexible image understanding across multiple domains and sensor types.
+IV-B Multi-level Prompting Data Annotation
+A meticulously designed multi-strategy annotation method is employed to label diverse multi-domain data, which is uniformly reorganized and formatted into image–visual prompt-text triples.
+Notably, visual prompts include bounding boxes, points, and free-form types.
+The annotation process is detailed as follows.
+IV-B1 Box prompts
+The regional data are based on detection datasets, and ground truth bounding boxes are used as visual prompts to recognize object-level categories.
+As shown in Fig.
+3 (a), an example of a SAR image is provided to illustrate the process.
+The traditional object detection data are transformed into multi-modal conversational samples, each comprising an image, visual prompts, and text (question-answer pairs).
+The image is directly sourced from the detection dataset, while the visual prompt is generated by applying slight perturbations to the ground-truth boxes, and the textual prompts are drawn from predefined templates.
+Specifically, image-level tasks use the boxes $0,0,\mathrm{width},\mathrm{height}]$, which can cover the entire image as the visual prompt.
+IV-B2 Point prompts
+To enable point-based interaction during training, we construct point prompt annotations by randomly sampling pixels within each ground-truth object mask from semantic segmentation datasets.
+As illustrated in Fig.
+3 (b), each image is divided into $32\times 32$ patches, and the points are randomly sampled within each patch as the visual prompts, where each point corresponds to a pixel-level label annotation.
+IV-B3 Free-form Prompts
+To enable the model to handle diverse and flexible visual inputs at inference time, we simulate free-form prompts through noise-based augmentation methods.
+Specifically, we consider both box and point prompts and augment them to reflect the imprecision and looseness of user-defined inputs, as illustrated in Fig.
+3 (c).
+First, for box prompt augmentation, to mimic diverse user-drawn free-form inputs, we apply Gaussian noise to the original ground-truth bounding boxes.
+Let $\mathcal{B}=(x,y,w,h)$ denote a bounding box parameterized by its top-left corner and dimensions.
+We generate an augmented box $\mathcal{B}^{\prime}$ via:
+$\mathcal{B}^{\prime}=\mathcal{B}+\mathcal{N}(0,\Sigma),\quad\text{where }\Sigma\propto\text{diag}(w^{2},h^{2},w^{2},h^{2}).$ (14)
+This noise injection perturbs both the location and scale of the box in a size-aware manner, leading to boxes that may loosely cover, shift from, or extend beyond the target object, thus statistically approximating free-form visual prompts.
+For point prompts, we simulate user interactions by randomly sampling multiple pixels within the ground-truth object mask and associating them with the same instance label.
+Given a binary mask $\mathcal{M}\subset\Omega$ representing a target object, we uniformly select $K$ points from the interior of $\mathcal{M}$.
+These sparse prompts guide the model to localize the target object via point-based reference.
+Additionally, at inference (see bottom part of Fig.
+3 (c)), free-form inputs are pre-processed into bounding boxes, enabling flexible, intuitive user-drawn prompts.
+IV-B4 Text Prompts
+To support instruction tuning on multi-source RS data, we provide a set of carefully designed natural language instructions that correspond to different visual prompts (e.g., box or point).
+Each instruction guides the model to perform a specific visual understanding task, such as classification, captioning, or relational reasoning.
+Designed task instructions are as follows:
+  • Please identify the object category of each marked region in the image.
+  • Please identify the labels of each marked point in the image.
+  • Please provide a brief caption of each marked region in the image.
+  • Please provide a detailed caption of each marked point in the image.
+  • Please analyze the relationship between all marked regions in the image.
+  • Please provide a summarized caption based on all the marked regions in the image.
+  • Please analyze how the marked objects interact with each other in the given scene.
+In addition, users can employ the <<Mark n>> or <<Region n>> identifiers to specify pixels or regions of interest and freely ask questions about them.
+These task-specific textual instructions guide the model to perform multiple visual tasks, promoting a more fine-grained understanding of visual content under different granularities.
+IV-C GPT4V-assisted Data Generation
+To enable high-quality annotations with GPT-4V [76], we followed the design principles of [77] and developed a set of rigorous and standardized prompt templates, as illustrated in Fig.
+4.
+These templates are tailored to distinct tasks: brief captioning, detailed captioning, and inter-relationship analysis.
+Notably, category names and ground-truth bounding boxes are explicitly embedded in the prompts, ensuring the correctness of the generated annotations by guiding GPT-4V to accurately associate text with visual regions.
+Beyond correctness, the structured templates also guarantee annotation consistency across large-scale data, avoiding ambiguity and hallucination.
+To further strengthen the coupling between visual prompts and textual outputs, we adopt Set-of-Marks (SoM) prompting [78], which highlights multiple target objects simultaneously and reinforces spatial-textual grounding.
+This design enables GPT-4V to produce coherent and complete descriptions even in cluttered scenes.
+In conclusion, the integration of diverse multi-domain datasets through this rigorous pipeline allows joint training of EarthGPT-X on standardized supervision, enhancing its robustness across RS modalities and improving its generalization to complex real-world scenarios.
+Additionally, to further ensure reliability, we conduct RS expert inspection on sampled annotations, validating semantic accuracy, coherence, and completeness.
+This human-in-the-loop verification step complements automatic generation and safeguards the overall quality of the constructed dataset.
 V Experiment
@@
 We use transformer-based decoder-only Llama 2-13B[79] as the basic LLM.
-The pre-trained DINOv2-ViT L/14 [56] and CLIP-ConvNeXt [55] are adopted as hybrid image encoder MVE.
+The pre-trained DINOv2-ViT L/14 [55] and CLIP-ConvNeXt [54] are adopted as a hybrid image encoder MVE.
 The training image resolution is $448\times 448$.
@@
 Throughout the training, to keep training efficiency and preserve the generalization ability of backbone models, we only unfreeze the self-attention layers of LLM and the added cross-attention module for tuning.
-Other modules including visual encoders, projection layers, etc., are kept frozen.
+Other modules, including visual encoders, projection layers, etc., are kept frozen.
 The AdamW optimizer[80] with a weight decay of 0.01 and a learning rate of 2e-5 is utilized.
-The total training is conducted on 8 NVIDIA A100 GPUs with 80 GB memory.
+The total training is conducted on 8 NVIDIA A100 GPUs with 80 GB memory, and the training time is around 310 hours.
 V-A2 Evaluation Metrics
-We adopt Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS)[73] as evaluation metrics on referring object classification tasks.
-Furthermore, we evaluate the performance of the proposed model on the region captioning tasks based on metrics for assessing language generation quality, including BLEU[74], METEOR[75], ROUGE[76], CIDER[77], and SPICE[78].
-Those metrics provide a diverse and comprehensive assessment of the quality of identified object information and generated region captions.
-TABLE II: Results on referring object classification of EarthGPT-X and other multiple MLLMs on DIOR-RSVG test set.
-Methods Publication Year Formats SS S-IoU
+We adopt three categories of evaluation metrics tailored to different visual tasks in our benchmark.
+(i) Region and Image Captioning tasks: We employ standard language generation metrics, including BLEU [81], METEOR [82], ROUGE [83], CIDEr [84], and SPICE [85].
+(ii) Scene Classification: We report classification accuracy, which directly reflects the model’s ability to recognize high-level scene semantics in RS images.
+(iii) Referring Object Classification: We employ two semantic-aware metrics: Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS) [86, 3, 77].
+SS measures the similarity of predicted and ground-truth labels in a semantic embedding space, while S-IOU captures the semantic overlap between them by computing the intersection-over-union in the embedding space.
+Those metrics provide a comprehensive assessment of the quality of the identified object information and generated captions.
+Consider the referring expression “A airplane is on the right of a small vehicle”, where the ground-truth target labels are airplane and vehicle.
+Suppose the model predicts aircraft and car, both outputs achieve high SS since they are semantically close to the labels in the embedding space, and also high S-IoU because of strong token-level overlap.
+In contrast, if the predictions are ship and building, SS drops due to weaker semantic similarity, while S-IoU becomes very low as the key object words differ significantly from the ground-truth labels.
+TABLE III: Zero-shot Scene Classification (UCMerced and AID).
+Models UCMerced Acc AID Acc
+Qwen-VL [87] 62.90 52.60
+MiniGPTV2 [88] 4.76 12.90
+LLaVa-1.5 [29] 68.00 51.00
+Sphinx [89] 62.76 58.20
+InternVL [90] 58.23 60.40
+GeoChat [9] 84.43 72.03
+EarthMarker [15] 86.52 77.97
+EarthGPT-X (Ours) 87.89 78.09
+TABLE IV: Referring object classification (DIOR-RSVG).
+Models Point Box (Coor) Free-form
+$S$ S-IoU $S$ S-IoU $S$ S-IoU
 Vanilla MLLMs
-GeoChat [9] CVPR 2024 Coor 79.59 68.80
-Sphinx [79] Arxiv 2023 Coor 93.72 89.37
-EarthGPT [8] TGRS 2024 Coor 94.64 90.16
-Natural Visual
-Prompting MLLMs
-Sphinx-V [50] Arxiv 2024 Box 89.07 81.62
-Vip-LLava-7b [80] CVPR 2024 Box 72.56 55.94
-Vip-LLava-13b [80] CVPR 2024 Box 74.51 60.53
-RS Visual
-Prompting MLLM
-EarthMarker [22] Arxiv 2024 Point 95.96 93.49
-EarthMarker [22] Arxiv 2024 Box 98.37 97.24
-EarthGPT-X(Ours) Point 96.02 94.12
-EarthGPT-X(Ours) Box 98.60 97.83
-EarthGPT-X(Ours) Free-form 98.30 96.56
-TABLE III: Results on region captioning of EarthGPT-X and other MLLMs on the OPT-RSVG test set.
-Method Publication Year Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDER SPICE
+InternVL2.5 [38] 54.54 37.91
+InternVL3 [91] 50.84 26.99
+GeoChat [9] 79.59 68.80
+Sphinx [89] 93.72 89.37
+EarthGPT [8] 94.64 90.16
+Visual Prompting MLLMs
+Sphinx -V [77] 86.21 78.78 89.07 81.62 88.46 79.97
+Vip-LLava-7b [18] 68.32 51.98 72.56 55.94 71.24 53.99
+Vip-LLava-13b [18] 70.02 57.61 74.51 60.53 74.05 59.83
+EarthMarker [15] 95.96 93.49 98.37 97.24
+EarthGPT-X (Ours) 96.02 94.12 98.60 97.83 98.30 96.56
+Notes: Since vanilla MLLMs lack visual prompt support, textual coordinates are treated as the referring regions. ‘—’ means no such capability.
+TABLE V: Referring Object Classification (MSAR).
+Models Formats $S$ S-IoU
+InternVL2.5 [38] Coor 37.05 7.69
+InternVL3 [91] Coor 32.72 2.89
+Sphinx-V [77] Box 34.76 2.03
+EarthGPT-X (Ours) Box 99.85 97.75
+TABLE VI: Supervised Image Captioning (NWPU-captions).
+Models BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGE-L CIDEr(0-5) SPICE
+Expert Models
+CSMLF [92] 77.0 64.9 53.2 47.1 32.0 57.8 106.5 26.5
+Qu, et. al [60] 72.5 60.3 51.8 45.5 33.6 59.1 117.9 27.6
+Attention (soft) [59] 73.1 60.9 52.5 46.2 33.9 59.9 113.6 28.5
+Attention (hard) [59] 73.3 61.0 52.7 46.4 34.0 60.0 110.3 28.4
+FC-Att [93] 73.6 61.5 53.2 46.9 33.8 60.0 123.1 28.3
+SM-Att [93] 73.9 61.7 53.2 46.8 33.0 59.3 123.6 27.6
+MLCA-Net [94] 74.5 62.4 54.1 47.8 33.7 60.1 126.4 28.5
+EarthGPT-X (Ours) 85.2 74.2 64.0 55.9 39.5 71.5 163.7 27.3
+Figure 5: Image brief and detailed captioning results on optical, infrared, and SAR modalities, demonstrating the superior scene-level visual understanding ability of EarthGPT-X.
+Figure 6: The referring object classification results on SAR and infrared RS images demonstrate the superior object-level RS visual understanding capability of EarthGPT-X, compared to other vanilla and visual prompting MLLMs (The black text indicates consistency with the ground truth, the red text represents incorrect answers).
+TABLE VII: Zero-shot region captioning (OPT-RSVG).
+Models Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr SPICE
 Vanilla MLLMs
-Qwen-VL-Chat [9] Arxiv 2023 Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
-GeoChat [81] CVPR 2024 Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
-Sphinx [79] Arxiv 2023 Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
-EarthGPT [8] TGRS 2024 Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
+Qwen-VL [9] Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
+InternVL2.5 [38] Coor 9.87 4.18 1.69 0.52 5.71 10.55 9.22 8.28
+InternVL3 [91] Coor 15.42 8.91 6.34 4.38 18.64 14.28 13.27 14.87
+GeoChat [87] Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
+Sphinx [89] Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
+EarthGPT [8] Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
 Visual Prompting MLLMs
-Sphinx-V [50] Arxiv 2024 Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
-Vip-LLava-7b [80] CVPR 2024 Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
-Vip-LLava-13b [80] CVPR 2024 Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
-GLaMM [80] CVPR 2024 Box 12.50 6.06 3.25 1.74 7.98 22.01 23.85 12.40
+Sphinx-V [77] Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
+Vip-LLava-7b [18] Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
+Vip-LLava-13b [18] Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
 EarthGPT-X (Ours) Point 52.22 48.13 45.15 42.88 35.01 64.31 477.40 63.14
@@
 EarthGPT-X (Ours) Free-form 63.25 58.16 54.89 52.01 38.35 65.01 482.26 63.23
-Figure 3: The referring object classification results on multi-source RS images demonstrate the superior region-level RS visual understanding capability of EarthGPT-X, compared to other vanilla and visual prompting MLLMs (the black text indicates consistency with the ground truth, the red text represents incorrect answers).
-Figure 4: Example visualization of EarthGPT-X for multi-source (optical, SAR, and infrared) image comprehension and multiple visual tasks reasoning under various visual prompts.
-V-B Referring Object Classification
-To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla MLLMs and visual prompting MLLMs on DIOR-RSVG[7].
-For referring object classification tasks, the text instruction is “Please identify the object category of the marked region in the image”.
+Note: The underline denotes the best zero-shot performance achieved under box prompts.
+TABLE VIII: Region Captioning on SAR and Infrared datasets.
+Models SAR-Ship (Supervised) HIT UAV (Zero-Shot)
+ROUGE-1 ROUGE-L METEOR ROUGE-1 ROUGE-L METEOR
+Generic MLLMs
+GPT-4o[95] 7.49 7.24 7.07 10.96 9.02 8.23
+InternVL2[90] 9.67 8.67 8.19 11.00 9.53 8.40
+Specialized RS MLLMs
+GeoChat[9] 57.15 57.15 52.20 59.85 59.85 51.31
+EarthDial[23] 63.10 63.10 54.83 61.83 61.83 52.80
+EarthGPT-X (Ours) 65.25 65.25 55.90 64.40 64.40 51.56
+Figure 7: Qualitative examples of EarthGPT-X performing multi-task reasoning and comprehending multi-source RS imagery (optical, SAR, and infrared) under various visual prompts.
+Figure 8: Illustration of multiple visual prompting strategies.
+Specifically, Point Prompt highlights local functional roles of individual regions, Box Prompt emphasizes global thematic coherence and spatial relations, and Free-form Prompt provides concise contextual abstraction.
+Despite these differences, all three prompting types guide the model to establish precise region–text alignment.
+TABLE IX: Ablation results about the design of the visual encoders.
+CNN Model ViT Model Classification Region Captioning Ref. Obj. Classification
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+V-B Scene Classification
+We evaluate EarthGPT-X on two RS benchmarks: AID [96] and UCMerced [97].
+Following GeoChat [9], we use 20% of AID for testing, while the entire UCMerced serves as a zero-shot test set.
+To fit the visual prompting setting, we define a full-image box $0,0,\mathrm{width},\mathrm{height}]$ and use the instruction: “Please identify the object category of each marked region in the image.” Classification accuracy is computed under zero-shot settings.
+As shown in Tab.
+III, EarthGPT-X achieves 78.09% on AID and 87.89% on UCMerced, outperforming all baselines.
+This further confirms that the proposed visual prompting paradigm achieves stronger task alignment than directly applying vanilla MLLMs (e.g., Sphinx and GeoChat) without prompt adaptation.
+V-C Referring Object Classification
+V-C1 Evaluation on Optical Data
+Referring object classification aims to identify the category labels of objects referred to by visual prompts and textual instructions.
+To assess the merits of the proposed EarthGPT-X, we report its performance and compare it with the recent vanilla and visual prompting MLLMs on DIOR-RSVG[7].
+We adapt DIOR-RSVG by converting referring expressions and their bounding boxes into classification samples, where each target region serves as input and its category label as output.
+The text instruction is “Please identify the object category of the marked region in the image”.
 The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions.
-The vanilla MLLMs including natural MLLMs Sphinx [79], RS MLLMs GeoChat [81], and EarthGPT [8].
-We select the most recent natural visual prompting MLLMs including Sphinx-V [50], Vip-LLava [80], and the RS visual prompting MLLM model EarthMarker[22] as the comparison.
-[…]
-Figure 5: Pixel grounding examples of EarthGPT-X on multi-source imagery (EarthGPT-X’s interpretation of the object classes is also displayed on the figures).
-V-C Region Captioning
-For the brief region captioning, the test set of OPT-RSVG [82] is employed for zero-shot evaluation.
+The vanilla MLLMs used for comparison include natural-domain models such as Sphinx [89], InternVL2.5-1B[38], InternVL3-1B[91], as well as RS-domain models including GeoChat [87] and EarthGPT [8].
+For visual prompting MLLMs, we select the most recent representatives, namely the natural-domain models Sphinx-V [77] and Vip-LLava [18], together with the RS-domain model EarthMarker [15].
+We conduct experiments under three levels of visual prompting (point, box, and free-form) to provide a fair and comprehensive comparison across models.
+Following [77, 1], for point prompts, we randomly select a point within the ground-truth object that is also close to its boundary.
+For box prompts, we use the ground-truth bounding box provided by DIOR-RSVG.
+For free-form prompts, we randomly perturb the center positions and zoom the box prompts to simulate.
+As shown in Tab.
+IV, EarthGPT-X achieves optimal results in S-IoU under box prompts.
+We also observe that the box-level results are better than the point and free-form prompt ones.
+V-C2 Evaluation on SAR Data
+We evaluate EarthGPT-X and recent InternVL series[38, 91], Sphinx-V[77] on the MSAR[69] test set (Tab.
+V).
+EarthGPT-X, trained with SAR-specific samples, achieves superior performance (SS: 99.85%, S-IoU: 97.75%), showing strong capacity in SAR-oriented understanding.
+In contrast, natural MLLMs InternVL2.5 and InternVL3, without SAR training, perform significantly lower.
+The superior results of EarthGPT-X illustrate its stronger modality adaptation and specialization capabilities, which are crucial for accurate SAR imagery understanding in the RS applications.
+V-D Image Captioning
+To evaluate the image captioning capabilities, we adopt the NWPU-Captions [94] dataset to benchmark EarthGPT-X against other expert models under supervised settings.
+During evaluation, we use the instruction “Please provide a brief caption of each marked region in the image” alongside a full-image box $0,0,\mathrm{width},\mathrm{height}]$ as the visual prompt to guide the captioning process.
+As reported in Tab.
+VI, EarthGPT-X consistently outperforms existing expert methods across all evaluation metrics.
+These results indicate that by treating the entire image as a unified region and leveraging multi-domain training on both natural and RS data, EarthGPT-X is capable of generating accurate and semantically rich descriptions for RS scenes.
+V-E Region Captioning
+V-E1 Evaluation on Optical Data
+For the brief region captioning on optical data, the test set of OPT-RSVG [98] is employed for zero-shot evaluation.
 We elaborately choose both natural and RS-specific MLLMs for performance comparison.
-Natural scene MLLMs contains Qwen-VL-Chat [9], Sphinx [79], Sphinx-V [50], Vip-LLava [80] and GLaMM [80].
+Natural scene MLLMs contains Qwen-VL [9], InternVL2.5-1B[38], InternVL3-1B[91], Sphinx [89], Sphinx-V [77], Vip-LLava [18] and GLaMM [18].
 Popular RS-specific MLLMs GeoChat [87] and EarthGPT [8] are selected.
 The text instruction for the region captioning task is designed as “Please provide a brief caption of each marked region in the image”.
+As shown in Tab.
+VII, the proposed EarthGPT-X consistently outperforms previous works across all evaluation metrics.
+Notably, EarthGPT-X achieves significant gains over Sphinx-V, despite the latter being trained on large-scale RS datasets.
+This highlights the superior performance and stronger generalization ability of EarthGPT-X in multi-source RS data understanding.
+Notably, using boxes as visual prompts can yield the best performance on region captioning ability.
+V-E2 Evaluation on SAR and Infrared Data
 As displayed in Tab.
-III, the proposed EarthGPT-X significantly outperforms the previous works on all of the evaluation metrics.
-Notably, using boxes as visual prompts can yield the best performance on region captioning and dialogue ability.
-V-D Comprehensive Capabilities Evaluation
+VIII, for the multi-source brief captioning task, we evaluate EarthGPT-X on two representative datasets: SAR-Ship [99] for supervised evaluation and HIT UAV [100] under a zero-shot setting.
+These datasets enable us to validate the generalization ability of EarthGPT-X across different modalities.
+For comparison, we include two representative general-purpose MLLMs, GPT-4o and InternVL2 [90].
+We further reference the first RS MLLM, GeoChat [9], and the latest multi-source understanding model, EarthDial [23], as strong domain-related baselines.
+Experimental results demonstrate that EarthGPT-X achieves superior overall performance on multi-source datasets.
+V-F Ablation Study
+TABLE X: Ablation results on Image Pre-processing.
+Settings Visual Encoder $S$ S-IoU
+Single-scale MVE 94.93 95.05
+Multi-scale MVE 96.18 95.29
+Multi-scale + Sub-images MVE 98.60 97.83
+TABLE XI: Computation Analysis and Comparisons
+ Models  Inference Time(s)  GPU Memory(MiB)
+MLLMs
+Sphinx 1.08 41853
+EarthGPT 0.79 41766
+GeoChat 0.74 15544
+Qwen-VL 0.35 38145
+Visual Prompting Models
+Sphinx-V 1.95 41794
+Vip-LLava-13B 0.84 27748
+EarthGPT-X(Ours) 1.60 41006
+V-F1 Different Visual Encoders
+To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders.
+Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison.
+The results are shown in Tab.
+IX, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieves the most superior performance on scene classification, region captioning, and referring object classification tasks.
+These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
+V-F2 Sub-images Decomposition and Multi-scale
+In the visual feature extraction stage, we strengthen feature integration by leveraging multi-scale features and high-resolution sub-images for richer RS image comprehension.
+Ablation results in Tab.
+X show that this design significantly outperforms single-scale settings on the DIOR-RSVG dataset, confirming that capturing finer details through multiple scales and sub-images is crucial for accurate interpretation.
+V-G Computation Analysis and Comparisons
+In this section, we compare the GPU memory usage (MiB) and inference time (seconds) of EarthGPT-X with other MLLMs and visual prompting models.
+All evaluations are conducted on the region captioning task using an NVIDIA RTX A6000.
+For both MLLMs and visual prompting models, GPU memory usage is primarily attributed to model weight storage.
+As illustrated in Tab.
+XI, EarthGPT-X consumes more GPU memory than Sphinx [77], GeoChat [9], EarthGPT [8], Qwen-VL [87], and Vip-LLava-13B [18], but less than Sphinx-V [77].
+Similarly, its inference time is longer than all other models except Sphinx-V, due to its larger parameter scale.
+However, the time difference remains within one second, indicating that the performance gap is marginal and does not affect deployment in practical applications, as EarthGPT-X remains well within the capacity of commonly used GPUs.
+In future work, we plan to optimize EarthGPT-X’s inference efficiency to better balance speed and interpretability.
+Figure 9: Examples of EarthGPT-X performing visual reasoning and pixel-level grounding on optical imagery across diverse scenes (harbor, urban stadium, and airport). Better to zoom in to view clearly.
+Figure 10: Examples of EarthGPT-X performing small vehicle recognition and reasoning with pixel-level grounding on multi-source imagery (i.e., SAR, infrared, optical–daytime, and optical–nighttime).
+Better to zoom in to view clearly.
+VI Visualization
+VI-A Multi-source Image Captioning
+We present visualization results of EarthGPT-X on optical, SAR, and infrared images to demonstrate its capability in understanding diverse RS scenes.
+Given an entire image as the region of interest, we use the visual prompt $0,0,\mathrm{width},\mathrm{height}]$ and provide instructions such as “Please provide a brief caption of each marked region in the image.
+” or “Please provide a detailed caption of each marked region in the image.” As shown in Fig.
+5, the model generates accurate and semantically rich captions across modalities, effectively handling varied visual characteristics and noise patterns.
+VI-B Referring Object Classification
+We conduct qualitative analysis to provide an intuitive performance display of EarthGPT-X.
+As shown in Fig.
+6, the infrared image showcases a residential leisure area, and for the three marked regions, all other MLLMs respond incorrectly.
+Conversely, EarthGPT-X successfully recognizes the human, vehicle, and basketball court targets.
+Following this is a fuzzy port SAR image, where it is difficult to visually distinguish between ships and storage tanks.
+EarthGPT-X delivers precise interpretations for these targets, whereas other models perform poorly, except for EarthGPT.
+These results prove the significance of our work that empowers the MLLMs with fine-grained reasoning ability for object-level comprehension.
+VI-C Multi-level Multi-task Reasoning
 Traditional vision models in the RS field primarily focus on a single visual modality and single tasks, exhibiting poor generalization[101, 102, 103].
@@
 As displayed in Fig.
-4, our model demonstrates remarkable visual reasoning and analytical abilities.
-EarthGPT-X can provide detailed descriptions for multi-source images, as shown in the interpretation results of the first image, which includes overall image depiction, number of targets, target orientation, and scene reasoning.
-Notably, in the relationship analysis of the second and third images, EarthGPT-X provides detailed and comprehensive responses.
-Specifically, EarthGPT-X adeptly summarizes the connections between various regions and the overall image relationship, and then analyzes the characteristics and attributes of the targets within each specific area.
+7, our model demonstrates remarkable visual reasoning and analytical abilities.
+EarthGPT-X provides detailed descriptions for multi-source images, as evidenced by the interpretation of the first image, which includes overall scene depiction, target counts, orientations, and higher-level reasoning.
+In the relationship analysis of the second and third examples, the model generates comprehensive and structured responses, effectively capturing both local and global semantics.
+Notably, even for low-quality SAR and infrared maritime images, EarthGPT-X can infer relative spatial relations, target sizes, vessel functionalities, and operational states (e.g., docking at a port).
 In summary, EarthGPT-X demonstrates a profound ability to dissect and understand complex visual information in the multi-source RS MLLMs.
-V-E Pixel Grounding
-To evaluate the pixel grounding capabilities of the proposed model, we specifically chose images that feature complex and different scenes with multiple targets.
-This experiment aims to test the model’s robustness and precision in identifying and segmenting diverse elements within a single frame.
+VI-D Multiple Visual Prompts
+To evaluate the adaptability of EarthGPT-X under different prompting shapes, we investigate point prompt, box prompt, and free-form prompt.
+These prompts vary in the way visual regions are highlighted to the model, thereby influencing the style and granularity of responses.
+i) Similarities: Across all prompt types, EarthGPT-X consistently grounds the responses to the marked regions, ensuring semantic coherence and correct alignment with the visual input.
+ii) Differences: As shown in the Fig.
+8, under the point prompt, EarthGPT-X generates a caption emphasizing the spatial relation, which reflects its local context-awareness.
+Under the box prompt, the model highlights the absolute spatial location, showing a region-level grounding behavior.
+In contrast, the free-form prompt yields a more generic description, focusing on object-level attributes (Note that these examples are qualitative illustrations, not unique or deterministic results, since different prompt placements may lead to varied outputs).
+VI-E Zero-shot Pixel-level Grounding and Reasoning
+All pixel grounding and reasoning experiments are conducted in a zero-shot setting, and the capability is obtained without any task-specific training.
+VI-E1 Different Scenes
+To evaluate the robustness under complex scenes, we select RS images with multiple diverse targets.
 As shown in Fig.
-5, the first test image is the port scene, we mark a ship and a tiny vehicle to assess the model’s ability to discern and accurately interpret tiny targets amidst a busy background.
-Surprisingly, EarthGPT-X not only recognized these smaller objects but also effectively segmented them from the surrounding environment.
-Another test involved an image from an urban sports center where we marked five distinct targets scattered across different areas, including varied elements such as a swimming pool, tennis court, vehicle, etc.
-This scenario tests the model’s capacity to handle spatially separated objects within a complex urban setting.
-EarthGPT-X demonstrates exceptional performances in precise object recognition and pixel-grounding.
-Moreover, these tests highlight EarthGPT-X’s advanced pixel understanding capabilities and its potential to enhance RS image analyses across different visual modalities.
-V-F Ablation Study
-In this section, we analyze the impact of key components of EarthGPT-X in detail.
-TABLE IV: Ablation results on different visual encoders.
-CLIP CLIP DINOv2 SS S-IoU
-ConvNeXt-L RN50$\times$16 ViT-L/14
-[…]
-[…]
-[…]
-[…]
-[…]
-TABLE V: Ablation results on Image Pre-processing.
-Settings Visual Encoder SS S-IoU
-Single-scale MVE 94.93 95.05
-Multi-scale MVE 96.18 95.29
-Multi-scale + Sub-images MVE 98.40 97.13
-V-F1 Different Visual Encoders
-To show the effectiveness of the proposed hybrid MVE module, we ablate on the performance of our EarthGPT-X when using different visual encoders.
-Specifically, this ablation uses CLIP-ConvNeXt-L, CLIP RN50×16, and DINOv2 ViT-L/14 for comparison.
-The results are shown in Tab.
-IV, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieve the most superior performance.
-The performance of both SS and S-IoU increases accordingly.
-These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
-V-F2 Sub-images Decomposition and Multi-scale
-In the visual feature extraction phase of our approach, we enhance the integration of visual features by incorporating multiple scales and high-resolution sub-images, enabling a more profound comprehension of RS images.
-To evaluate the effectiveness of our modified feature extraction strategy, we conduct ablation studies focusing on the impact of multi-scale features and high-resolution sub-images on the performance of our EarthGPT-X model.
-As shown in Tab.
-V, utilizing the multi-scale and sub-images features setting achieves superior performance over single-scale features.
-These results demonstrate that the detailed analysis of these high-resolution sub-images and more scales allows for capturing finer details that are crucial for accurate interpretation and analysis.
-VI Conclusion
+9, in the port scene, we mark a ship and a tiny vehicle to examine the model’s ability to localize small objects within a cluttered background.
+EarthGPT-X not only successfully detects these targets but also separates them from the surroundings.
+In another urban sports center scene, five spatially separated objects (e.g., swimming pool, tennis court, and vehicle) are interpreted.
+EarthGPT-X demonstrates exceptional performance in precise object recognition and pixel-grounding.
+VI-E2 Different Image Modalities
+As shown in Fig.
+10, we further evaluate on small vehicle targets across heterogeneous image modalities, including optical, infrared, and SAR.
+The test data is from VisDrone-DroneVehicle [104] and SAR_Vehicle_Detection_Dataset[105].
+Despite the significant domain gaps among these modalities, EarthGPT-X consistently identifies and segments small and tiny vehicles, showing its strong generalization and adaptability to heterogeneous RS image types.
+VII Conclusion
 In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources.
-As the first RS spatial model to unify referring and grounding tasks within a visual prompting framework, EarthGPT-X represents a significant advancement in the application of MLLMs to the RS domain.
 Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery.
-Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios, such as SAR and infrared.
-Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-grained visual understanding at the image, region, pixel levels, etc., allowing for precise and intelligent analyses in real-world applications.
-In the future, we plan to enrich the SAR and infrared datasets to encompass a broader range of tasks, addressing the current limitation of narrow scope.
-Additionally, we aim to enable EarthGPT-X to perform more complex visual reasoning tasks across multiple visual modalities, further solidifying EarthGPT-X as a versatile and powerful tool for comprehensive RS imagery analyses.
+Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios.
+Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-level visual understanding at the image, region, and pixel levels, etc., allowing for precise and intelligent analyses in real-world applications.
+In the future, we aim to expand EarthGPT-X towards a wider range of tasks and more advanced visual reasoning capabilities across diverse modalities, ultimately strengthening its versatility and effectiveness in comprehensive RS imagery analysis.

### v3

@@
 Distinct from prior models, EarthGPT-X introduces: 1) a dual-prompt mechanism combining text instructions with various visual prompts (i.e., point, box, and free-form) to mimic the versatility of referring in human life; 2) a comprehensive multi-source multi-level prompting dataset, the model advances beyond holistic image understanding to support hierarchical spatial reasoning, including scene-level understanding and fine-grained object attributes and relational analysis; 3) a cross-domain one-stage fusion training strategy, enabling efficient and consistent alignment across modalities and tasks.
-Extensive experiments demonstrate that EarthGPT-X substantially outperforms prior nature and RS MLLMs, establishing the first framework capable of multi-source, multi-task, and multi-level interpretation using visual prompting in RS scenarios.
+Extensive experiments demonstrate that EarthGPT-X substantially outperforms prior natural and RS MLLMs, establishing the first framework capable of multi-source, multi-task, and multi-level interpretation using visual prompting in RS scenarios.
 The code and dataset are available at https://github.com/wivizhang/EarthGPT-X.
@@
 Typical spatial understanding tasks in multi-modal learning require aligning image regions with text semantics[1].
-Beyond holistic description, such tasks involve accurately interpreting specific regions or objects based on user instructions[2, 3].
+Beyond holistic analysis, spatial tasks involve accurately interpreting specific regions or objects based on user instructions[2, 3].
 Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focus of research in natural scenarios[4, 5, 6].
@@
 Additionally, Wu et al.
-[13] provide a review of visual prompting approaches, including visual prompting, prompt generation, and prompt learning.
-However, designed for natural images, these models cannot be directly adapted to the heterogeneous RS domain due to fundamental differences in sensing physics, modalities, and scales.
-Figure 1: Overview of the proposed framework for flexible and comprehensive spatial understanding in RS imagery: (a) Tasks: a wide range of multi-granularity visual tasks; (b) Model: EarthGPT-X integrates an image encoder, a visual prompt encoder, a text tokenizer, an LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is constructed from multi-source RS data (optical, SAR, infrared) and further enriched with GPT-4V–based annotations.
+[12] provide a review of visual prompting approaches, including prompt generation and prompt learning.
+However, designed for natural images, these models cannot be directly adapted to the heterogeneous RS domain due to fundamental differences in sensing physics, modalities, and spatial resolutions.
+Figure 1: Overview of the proposed framework for flexible and comprehensive spatial understanding in RS imagery: (a) Tasks: a wide range of multi-level visual reasoning tasks; (b) Model: EarthGPT-X integrates an image encoder, a visual prompt encoder, a text tokenizer, an LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is constructed from multi-source RS data (i.e., optical, synthetic aperture radar (SAR), infrared) and further enriched with GPT-4V–based annotations.
 Limited efforts in the RS domain have been made to enhance spatial understanding.
@@
 Thus, enabling MLLMs to interpret multi-source RS imagery at any granularity remains in its infancy.
-To overcome these limitations, we propose EarthGPT-X, a spatial MLLM built upon visual prompt learning, tailored for multi-source RS imagery, supporting diverse visual prompt types (e.g., points, boxes, circles, and scribbles), and enabling comprehensive interpretation from overall scene understanding to fine-grained regional analysis, thereby offering more flexible and scalable visual-language interaction than prior approaches.
+To overcome these limitations, we propose EarthGPT-X, a spatial MLLM built upon visual prompt learning and tailored for multi-source RS imagery.
+It supports arbitrary visual prompt types (e.g., points, boxes, circles, and scribbles) and enables comprehensive interpretation from overall scene understanding to fine-grained regional analysis, thereby providing more flexible and scalable visual–language interaction than prior approaches.
 As depicted in Fig.
-1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, and a construction pipeline of training datasets.
-Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification).
+1, the EarthGPT-X architecture consists of three interconnected components: multi-level RS image understanding tasks, a visual prompting model, and a construction pipeline of training dataset.
+Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analysis), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification).
 Secondly, the key of the model is to realize multi-modal content integration, including images, visual prompts, and text instructions.
@@
 The results indicate that EarthGPT-X achieves accurate multi-source reasoning and provides more detailed RS image understanding.
-Overall, EarthGPT-X not only attains competitive performance on vision–language benchmarks, but also demonstrates remarkable advantages in complex object-level reasoning tasks.
+Overall, EarthGPT-X not only attains competitive performance on visual–language benchmarks, but also demonstrates remarkable advantages in complex object-level reasoning tasks.
 Our contributions can be summarized as follows.
@@
 The first multi-source visual prompting dataset for RS, titled M-RSVP, consisting of over 650k image-visual prompt-text triples, is constructed.
-The constructed M-RSVP bridges the gap of multi-source visual prompting data in the RS domain, enabling multi-granularity interpretation across diverse scenes and sensors.
+The constructed M-RSVP bridges the gap of multi-source visual prompting data in the RS domain, enabling multi-granularity image interpretation across diverse scenes and sensors.
   • Multi-domain Multi-modal Multi-level Interaction.
 A hybrid signal mutual understanding mechanism is designed to strengthen the interplay between dense image features, sparse visual prompts, and text instructions, thereby improving the interpretation of referring regions within holistic images.
-Moreover, our cross-domain one-stage fusion training significantly streamlines the training procedure of MLLMs, while ensuring seamless integration of multi-source knowledge.
+Moreover, our cross-domain one-stage fusion training significantly streamlines the training procedure of MLLMs, while ensuring seamless integration of multi-modal knowledge.
   • Superior Performance.
-Extensive experiments benchmark EarthGPT-X against SOTA natural and RS MLLMs across diverse multi-level visual tasks and multi-source imagery.
+Extensive experiments benchmark EarthGPT-X against SOTA natural and RS MLLMs across diverse multi-level visual tasks and multi-source imagery, outperforming existing models.
 Furthermore, EarthGPT-X also excels in complex reasoning tasks, such as object inter-relationship analysis.
@@
 II-C Prompting Learning
-Prompt learning has become a key research direction in NLP [24], with works such as AutoPrompt [50] and CoOp [51] automating prompt design for language and vision–language models.
+Prompt learning has become a key research direction in NLP [23], with works such as AutoPrompt [49] and CoOp [50] automating prompt design for language and visual–language models.
 Subsequently, the idea evolved into visual prompting [12], marking a broader shift from language-based to multi-modal prompts.
-A milestone in this line is the Segment Anything Model (SAM) [12], which enables class-agnostic segmentation via point, box, or text prompts and demonstrates strong generalization across diverse domains.
+A milestone in this line is the SAM [11], which enables class-agnostic segmentation via point, box, or text prompts and demonstrates strong generalization across diverse domains.
 GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition.
@@
 It integrates images, visual prompts, and text instructions through a hybrid signal interaction mechanism, and employs a unified one-stage fusion training strategy to enable multi-source knowledge integration and multi-level multi-task reasoning.
-The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module.
-The multi-modal content encoding and integration are detailed as follows.
+The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure. The multi-modal content encoding and integration are detailed as follows.
 III-A1 Visual Global and Local Perception
@@
 $H_{\mathrm{ca}}$ $=\mathrm{Cross\text{-}attention}~(V_{\mathrm{sa}},~E_{\mathrm{ff}}),$ (6)
-$[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~(V_{\mathrm{ca}}+R_{\mathrm{ff}}),$ (7)
+$[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~(H_{\mathrm{ca}}+E_{\mathrm{ff}}),$ (7)
 where $V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts.
@@
 The binary segmentation mask $M$ generated process can be represented as follows
-$M=\mathcal{P}\Big(\mathcal{R}_{\mathrm{lp}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big).$ (13)
+$M=\mathcal{P}\Big(\mathcal{R}_{\mathrm{lp\text{-}p}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big).$ (13)
 In summary, EarthGPT-X, built on a tripartite architecture and an all-in-one training strategy, enables effective cross-modal and multi-source knowledge comprehension.
@@
 IV-C GPT4V-assisted Data Generation
-To enable high-quality annotations with GPT-4V [76], we followed the design principles of [77] and developed a set of rigorous and standardized prompt templates, as illustrated in Fig.
+To enable high-quality annotations with GPT-4V [75], we follows the design principles of [76] and developed a set of rigorous and standardized prompt templates, as illustrated in Fig.
 4.
@@
 SS measures the similarity of predicted and ground-truth labels in a semantic embedding space, while S-IOU captures the semantic overlap between them by computing the intersection-over-union in the embedding space.
-Those metrics provide a comprehensive assessment of the quality of the identified object information and generated captions.
-Consider the referring expression “A airplane is on the right of a small vehicle”, where the ground-truth target labels are airplane and vehicle.
+These metrics provide a comprehensive assessment of the quality of the identified object information and generated captions.
+Consider the referring expression “An airplane is on the right of a small vehicle”, where the ground-truth target labels are airplane and vehicle.
 Suppose the model predicts aircraft and car, both outputs achieve high SS since they are semantically close to the labels in the embedding space, and also high S-IoU because of strong token-level overlap.
 In contrast, if the predictions are ship and building, SS drops due to weaker semantic similarity, while S-IoU becomes very low as the key object words differ significantly from the ground-truth labels.
+V-B Scene Classification
+We evaluate EarthGPT-X on two RS benchmarks: AID [86] and UCMerced [87].
+Following GeoChat [8], we use 20% of AID for testing, while the entire UCMerced serves as a zero-shot test set.
+To fit the visual prompting setting, we define a full-image box $0,0,\mathrm{width},\mathrm{height}]$ and use the instruction: “Please identify the object category of each marked region in the image.” Classification accuracy is computed under zero-shot settings.
+As shown in Tab.
+III, EarthGPT-X achieves 78.09% on AID and 87.89% on UCMerced, outperforming all baselines.
+This further confirms that the proposed visual prompting paradigm achieves stronger task alignment than directly applying vanilla MLLMs (e.g., Sphinx and GeoChat) without prompt adaptation.
 TABLE III: Zero-shot Scene Classification (UCMerced and AID).
-Models UCMerced Acc AID Acc
-Qwen-VL [87] 62.90 52.60
-MiniGPTV2 [88] 4.76 12.90
-LLaVa-1.5 [29] 68.00 51.00
-Sphinx [89] 62.76 58.20
-InternVL [90] 58.23 60.40
-GeoChat [9] 84.43 72.03
-EarthMarker [15] 86.52 77.97
+Models Publications UCMerced Acc AID Acc
+Qwen-VL [88] arXiv 2023 62.90 52.60
+MiniGPTV2 [89] arXiv 2023 4.76 12.90
+LLaVa-1.5 [28] NeurIPS 2023 68.00 51.00
+Sphinx [90] ECCV 2024 62.76 58.20
+InternVL [91] CVPR 2024 58.23 60.40
+GeoChat [8] CVPR 2024 84.43 72.03
+EarthMarker [14] TGRS 2025 86.52 77.97
 EarthGPT-X (Ours) 87.89 78.09
 TABLE IV: Referring object classification (DIOR-RSVG).
-Models Point Box (Coor) Free-form
+Models Publications Point Box (Coor) Free-form
 $S$ S-IoU $S$ S-IoU $S$ S-IoU
 Vanilla MLLMs
-InternVL2.5 [38] 54.54 37.91
-InternVL3 [91] 50.84 26.99
-GeoChat [9] 79.59 68.80
-Sphinx [89] 93.72 89.37
-EarthGPT [8] 94.64 90.16
+InternVL2.5 [37] arXiv 2024 54.54 37.91
+InternVL3 [92] arXiv 2025 50.84 26.99
+GeoChat [8] CVPR 2024 79.59 68.80
+Sphinx [90] ECCV 2024 93.72 89.37
+EarthGPT [7] TGRS 2024 94.64 90.16
 Visual Prompting MLLMs
-Sphinx -V [77] 86.21 78.78 89.07 81.62 88.46 79.97
-Vip-LLava-7b [18] 68.32 51.98 72.56 55.94 71.24 53.99
-Vip-LLava-13b [18] 70.02 57.61 74.51 60.53 74.05 59.83
-EarthMarker [15] 95.96 93.49 98.37 97.24
-EarthGPT-X (Ours) 96.02 94.12 98.60 97.83 98.30 96.56
-Notes: Since vanilla MLLMs lack visual prompt support, textual coordinates are treated as the referring regions. ‘—’ means no such capability.
+Vip-LLava-7b [17] CVPR 2024 68.32 51.98 72.56 55.94 71.24 53.99
+Vip-LLava-13b [17] CVPR 2024 70.02 57.61 74.51 60.53 74.05 59.83
+Sphinx -V [76] ICLR 2025 86.21 78.78 89.07 81.62 88.46 79.97
+EarthMarker [14] TGRS 2025 95.96 93.49 98.37 97.24
+EarthGPT-X (Ours) 96.33 94.42 98.79 98.03 98.71 96.72
+Notes: Since vanilla MLLMs lack visual prompt support, textual coordinates (abbr. as Coor) are treated as the referring regions. ‘—’ means no such capability.
+The underline denotes the best performance achieved under box prompts.
 TABLE V: Referring Object Classification (MSAR).
-Models Formats $S$ S-IoU
-InternVL2.5 [38] Coor 37.05 7.69
-InternVL3 [91] Coor 32.72 2.89
-Sphinx-V [77] Box 34.76 2.03
+Models Publications Formats $S$ S-IoU
+InternVL2.5 [37] arXiv 2024 Coor 37.05 7.69
+InternVL3 [92] arXiv 2025 Coor 32.72 2.89
+Sphinx-V [76] ICLR 2025 Box 34.76 2.03
 EarthGPT-X (Ours) Box 99.85 97.75
 TABLE VI: Supervised Image Captioning (NWPU-captions).
-Models BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGE-L CIDEr(0-5) SPICE
-Expert Models
-CSMLF [92] 77.0 64.9 53.2 47.1 32.0 57.8 106.5 26.5
-Qu, et. al [60] 72.5 60.3 51.8 45.5 33.6 59.1 117.9 27.6
-Attention (soft) [59] 73.1 60.9 52.5 46.2 33.9 59.9 113.6 28.5
-Attention (hard) [59] 73.3 61.0 52.7 46.4 34.0 60.0 110.3 28.4
-FC-Att [93] 73.6 61.5 53.2 46.9 33.8 60.0 123.1 28.3
-SM-Att [93] 73.9 61.7 53.2 46.8 33.0 59.3 123.6 27.6
-MLCA-Net [94] 74.5 62.4 54.1 47.8 33.7 60.1 126.4 28.5
+Models Publications BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE -L CIDEr(0-5) SPICE
+RS Expert Models
+CSMLF [93] GRSL 2019 77.0 64.9 53.2 47.1 32.0 57.8 106.5 26.5
+Qu et al. [59] CITS 2016 72.5 60.3 51.8 45.5 33.6 59.1 117.9 27.6
+Attention (soft) [58] TGRS 2017 73.1 60.9 52.5 46.2 33.9 59.9 113.6 28.5
+Attention (hard) [58] TGRS 2017 73.3 61.0 52.7 46.4 34.0 60.0 110.3 28.4
+FC-Att [94] TGRS 2019 73.6 61.5 53.2 46.9 33.8 60.0 123.1 28.3
+SM-Att [94] TGRS 2019 73.9 61.7 53.2 46.8 33.0 59.3 123.6 27.6
+MLCA-Net [95] TGRS 2022 74.5 62.4 54.1 47.8 33.7 60.1 126.4 28.5
+Capformer [96] IGARSS 2022 89.0 81.1 74.5 69.2 45.4 78.4 201.9
+Aware-Transformer [97] CGI 2023 91.5 85.2 79.6 75.0 47.7 81.1 214.7
+RS MLLMs
+RS-CapRet [98] arXiv 2024 87.1 78.7 71.7 65.6 43.6 77.6 192.9 31.1
+EarthGPT [7] TGRS 2024 87.1 78.7 71.6 65.5 44.5 78.2 192.6 32.2
+EarthMarker [14] TGRS 2025 84.4 73.1 62.9 54.3 37.5 70.0 162.9 26.8
 EarthGPT-X (Ours) 85.2 74.2 64.0 55.9 39.5 71.5 163.7 27.3
@@
 TABLE VII: Zero-shot region captioning (OPT-RSVG).
-Models Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr SPICE
+Models Publications Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr SPICE
 Vanilla MLLMs
-Qwen-VL [9] Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
-InternVL2.5 [38] Coor 9.87 4.18 1.69 0.52 5.71 10.55 9.22 8.28
-InternVL3 [91] Coor 15.42 8.91 6.34 4.38 18.64 14.28 13.27 14.87
-GeoChat [87] Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
-Sphinx [89] Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
-EarthGPT [8] Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
+Qwen-VL [88] arXiv 2023 Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
+InternVL2.5 [37] arXiv 2024 Coor 9.87 4.18 1.69 0.52 5.71 10.55 9.22 8.28
+InternVL3 [92] arXiv 2025 Coor 15.42 8.91 6.34 4.38 18.64 14.28 13.27 14.87
+GeoChat [8] CVPR 2024 Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
+Sphinx [90] ECCV 2024 Box 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
+EarthGPT [7] TGRS 2024 Box 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
 Visual Prompting MLLMs
-Sphinx-V [77] Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
-Vip-LLava-7b [18] Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
-Vip-LLava-13b [18] Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
-EarthGPT-X (Ours) Point 52.22 48.13 45.15 42.88 35.01 64.31 477.40 63.14
-EarthGPT-X (Ours) Box 64.55 59.76 56.10 53.11 39.22 66.64 487.44 64.20
-EarthGPT-X (Ours) Free-form 63.25 58.16 54.89 52.01 38.35 65.01 482.26 63.23
+Sphinx-V [76] ICLR 2025 Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
+Vip-LLava-7b [17] CVPR 2024 Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
+Vip-LLava-13b [17] CVPR 2024 Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
+GLaMM [99] CVPR 2024 Box 12.50 6.06 3.25 1.74 7.98 22.01 23.85 12.40
+EarthGPT-X (Ours) Point 52.57 48.14 45.72 43.38 35.28 64.63 479.92 63.47
+EarthGPT-X (Ours) Box 64.83 59.91 56.28 53.64 39.61 66.87 488.76 64.32
+EarthGPT-X (Ours) Free-form 63.38 58.62 55.12 52.41 38.72 65.43 483.63 63.58
 Note: The underline denotes the best zero-shot performance achieved under box prompts.
 TABLE VIII: Region Captioning on SAR and Infrared datasets.
-Models SAR-Ship (Supervised) HIT UAV (Zero-Shot)
+Models Publications SAR-Ship (Supervised) HIT UAV (Zero-Shot)
 ROUGE-1 ROUGE-L METEOR ROUGE-1 ROUGE-L METEOR
 Generic MLLMs
-GPT-4o[95] 7.49 7.24 7.07 10.96 9.02 8.23
-InternVL2[90] 9.67 8.67 8.19 11.00 9.53 8.40
+GPT-4o [100] OpenAI 2024 7.49 7.24 7.07 10.96 9.02 8.23
+InternVL2 [91] OpenGV 2024 9.67 8.67 8.19 11.00 9.53 8.40
 Specialized RS MLLMs
-GeoChat[9] 57.15 57.15 52.20 59.85 59.85 51.31
-EarthDial[23] 63.10 63.10 54.83 61.83 61.83 52.80
+GeoChat [8] CVPR 2024 57.15 57.15 52.20 59.85 59.85 51.31
+EarthDial [22] CVPR 2025 63.10 63.10 54.83 61.83 61.83 52.80
 EarthGPT-X (Ours) 65.25 65.25 55.90 64.40 64.40 51.56
@@
 […]
-V-B Scene Classification
-We evaluate EarthGPT-X on two RS benchmarks: AID [96] and UCMerced [97].
-Following GeoChat [9], we use 20% of AID for testing, while the entire UCMerced serves as a zero-shot test set.
-To fit the visual prompting setting, we define a full-image box $0,0,\mathrm{width},\mathrm{height}]$ and use the instruction: “Please identify the object category of each marked region in the image.” Classification accuracy is computed under zero-shot settings.
-As shown in Tab.
-III, EarthGPT-X achieves 78.09% on AID and 87.89% on UCMerced, outperforming all baselines.
-This further confirms that the proposed visual prompting paradigm achieves stronger task alignment than directly applying vanilla MLLMs (e.g., Sphinx and GeoChat) without prompt adaptation.
+Notes: Results for region-level tasks are reported under box-prompt settings.
 V-C Referring Object Classification
@@
 The vanilla MLLMs do not support drawing visual prompts on the images, therefore, the coordinates contained in the text instruction serve as the referring regions.
-The vanilla MLLMs used for comparison include natural-domain models such as Sphinx [89], InternVL2.5-1B[38], InternVL3-1B[91], as well as RS-domain models including GeoChat [87] and EarthGPT [8].
+The vanilla MLLMs used for comparison include natural-domain models such as Sphinx [90], InternVL2.5[37], InternVL3[92], as well as RS-domain models including GeoChat [88] and EarthGPT [7].
 For visual prompting MLLMs, we select the most recent representatives, namely the natural-domain models Sphinx-V [76] and Vip-LLava [17], together with the RS-domain model EarthMarker [14].
@@
 During evaluation, we use the instruction “Please provide a brief caption of each marked region in the image” alongside a full-image box $0,0,\mathrm{width},\mathrm{height}]$ as the visual prompt to guide the captioning process.
-As reported in Tab.
-VI, EarthGPT-X consistently outperforms existing expert methods across all evaluation metrics.
-These results indicate that by treating the entire image as a unified region and leveraging multi-domain training on both natural and RS data, EarthGPT-X is capable of generating accurate and semantically rich descriptions for RS scenes.
+As shown in Table VI, EarthGPT-X outperforms most expert-designed models but remains inferior to task-specific architectures such as Aware-Transformer [97].
+We note that the performance gap in image captioning is a common limitation of MLLMs, which aim to balance multiple capabilities rather than optimizing for a single task.
 V-E Region Captioning
@@
 We elaborately choose both natural and RS-specific MLLMs for performance comparison.
-Natural scene MLLMs contains Qwen-VL [9], InternVL2.5-1B[38], InternVL3-1B[91], Sphinx [89], Sphinx-V [77], Vip-LLava [18] and GLaMM [18].
+Natural scene MLLMs contain Qwen-VL [8], InternVL2.5[37], InternVL3[92], Sphinx [90], Sphinx-V [76], Vip-LLava [17] and GLaMM [17].
 Popular RS-specific MLLMs GeoChat [88] and EarthGPT [7] are selected.
@@
 These datasets enable us to validate the generalization ability of EarthGPT-X across different modalities.
-For comparison, we include two representative general-purpose MLLMs, GPT-4o and InternVL2 [90].
+For comparison, we include two representative general-purpose MLLMs, GPT-4o from OpenAI and InternVL2 [91] from OpenGVLab.
 We further reference the first RS MLLM, GeoChat [8], and the latest multi-source understanding model, EarthDial [22], as strong domain-related baselines.
@@
 Multi-scale MVE 96.18 95.29
-Multi-scale + Sub-images MVE 98.60 97.83
+Multi-scale + Sub-images MVE 98.79 98.03
+Notes: Results are reported under box-prompt settings.
 TABLE XI: Computation Analysis and Comparisons
@@
 The results are shown in Tab.
-IX, it is clear that the setting using the ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieves the most superior performance on scene classification, region captioning, and referring object classification tasks.
+IX, it is clear that the setting using the CLIP-ConvNeXt-L and DINOv2 ViT-L/14 as the mixed visual experts achieves the most superior performance on scene classification, region captioning, and referring object classification tasks.
 These results demonstrate that utilizing a network architecture with complementary functions has deepened the understanding of local information and contextual dependencies on RS imagery.
@@
 Ablation results in Tab.
-X show that this design significantly outperforms single-scale settings on the DIOR-RSVG dataset, confirming that capturing finer details through multiple scales and sub-images is crucial for accurate interpretation.
+X show that this design significantly outperforms single-scale settings under the referring object classification task, confirming that capturing finer details through multiple scales and sub-images is crucial for accurate interpretation.
 V-G Computation Analysis and Comparisons
@@
 i) Similarities: Across all prompt types, EarthGPT-X consistently grounds the responses to the marked regions, ensuring semantic coherence and correct alignment with the visual input.
-ii) Differences: As shown in the Fig.
+ii) Differences: As shown in Fig.
 8, under the point prompt, EarthGPT-X generates a caption emphasizing the spatial relation, which reflects its local context-awareness.
@@
 Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-level visual understanding at the image, region, and pixel levels, etc., allowing for precise and intelligent analyses in real-world applications.
-In the future, we aim to expand EarthGPT-X towards a wider range of tasks and more advanced visual reasoning capabilities across diverse modalities, ultimately strengthening its versatility and effectiveness in comprehensive RS imagery analysis.
+In the future, we plan to enhance the image captioning capability of EarthGPT-X through task-aware fine-tuning and adaptive prompt optimization.
+Furthermore, we aim to expand EarthGPT-X towards a wider range of tasks and more advanced visual reasoning capabilities across diverse modalities, ultimately strengthening its versatility and effectiveness in comprehensive RS imagery analysis.